### v3

# MSCCL++: Rethinking GPU Communication Abstractions for Cutting-Edge AI Applications

Abstract.
[…]
conference: ; ; Now at Microsoft AI. Now at xAI.
1. Introduction
GPU communication has become a key area of optimization for high-performance AI applications. For instance, modern Large Language Models (LLMs) (Dubey et al., 2024; Brown et al., 2020) are often designed to run a single inference task on tens of GPUs (spanning a few physical nodes) to distribute computation and minimize end-to-end latency. Training tasks often scale much further up to tens of thousands of GPUs (Dubey et al., 2024). GPUs are connected and communicate with each other through either intra-node (PCIe, NVLink (NVIDIA, 2025d), xGMI (AMD, 2023), etc.) or inter-node (Ethernet or InfiniBand (IB) (inf, 2023)) links. Efficient communication is challenging due to the collective communication patterns of AI workloads that frequently distribute and gather data across all GPUs simultaneously. For example, communication patterns such as AllReduce, AllGather, AllToAll, Broadcast, and ReduceScatter (Wikipedia, 2025) are frequently used in modern AI applications to distribute computing workload over many GPUs and collect computed results into a single GPU.
A substantial portion of the end-to-end latency is often attributed to GPU communication in real workloads, usually in the 10% - 40% range of end-to-end LLM workloads. For example, a GPT-3 (Brown et al., 2020) model inference may spend 30% of the end-to-end time in AllReduce communication. Many efforts have been made to optimize GPU communication, such as developing efficient communication algorithms (routing paths and transfer scheduling) (Shah et al., 2023; Cao et al., 2024; Liu et al., 2024), overlapping communication with computation (Jangda et al., 2022; Wang et al., 2023), and other lower-level optimizations in the stack (NVIDIA, 2025a; AMD, 2025).
Despite these efforts, achieving high communication performance is still challenging and time consuming in real applications. Specifically, practitioners still need to write custom communication code to achieve the best performance, often from scratch. For instance, TensorRT-LLM (NVIDIA, 2025e), a popular LLM framework featured by NVIDIA in recent years, implements custom AllReduce communication methods from scratch. This is notable given that there already exists a popular GPU communication library developed by NVIDIA for many years, the NVIDIA Collective Communication Library (NCCL) (NVIDIA, 2025a). The custom communication of TensorRT-LLM outperforms NCCL in a wide range of LLM scenarios, especially when the data size is relatively small, while TensorRT-LLM still uses NCCL for larger data sizes. This raises a simple question: Why build a new ad hoc stack from scratch, instead of developing on top of an existing stack?
Surprisingly, such ad hoc software development has already become a widely common practice in cutting-edge AI applications. From our empirical studies, we find that the root cause lies in the fast evolving hardware. The enormous computing demand by modern AI applications is pushing the industry to aggressively upgrade chips and interconnects, resulting in powerful but immature hardware that is dramatically different from the previous generation. Cutting-edge AI applications are usually targeted for deployment over the latest hardware to obtain the best performance and efficiency. Thus, it is not surprising for existing general-purpose libraries (such as NCCL) to perform sub-optimally for those applications because it takes time for those libraries to optimize for diverse workload scenarios. Consequently, production implementations go with custom quick-and-dirty optimization specific to their own workloads and hardware environments. This incurs redundant development efforts across applications and writes a lot of non-portable code, which leads to further development effort when a new hardware becomes available.
In this paper, we discuss an alternative interface design for AI software libraries to support portability and custom optimizations by reducing redundant efforts without harming the flexibility of customization. The key idea is to separate high-level abstractions and optimizations from primitive hardware abstractions based on two observations. First, existing hardware abstractions like NCCL primitives try to hide hardware features to make it easier to program, but may hinder optimizations. Instead, a low-level, primitive interface that minimally abstracts the hardware would make it easier for: (1) application developers to optimize and fine-tune by enabling precise control of the hardware, and (2) library developers to support new hardware features by only providing a shallow layer of abstraction over it. Second, existing collective communication libraries (like NCCL) and DSLs (domain-specific languages; like MSCCLang (Cowan et al., 2023)) provide ease of use, but may lack workload-specific optimizations. Such higher level abstractions could provide portability and enable users to trade-off performance for programming effort.
Based on these insights, we propose MSCCL++ (Microsoft Collective Communication Library ++, pronounced em-sickle plus-plus), a novel GPU communication stack designed for high-performance AI applications. Unlike existing libraries, MSCCL++ exposes the most fundamental building blocks of communication as a user interface, namely the primitive interface (MSCCL++ Primitive API). This interface is very close to GPU hardware as a shallow layer over low-level GPU instructions, thus enabling quick adoption of latest hardware features with low effort. The primitive interface abstracts straightforward concepts of communication such as put, signal, wait, and flush. This helps translate low-level hardware behaviors into higher-level insights (e.g., bandwidth and latency) for optimizing communication. The primitive interface is designed to be very flexible to enable various optimizations that have been difficult to implement with existing stacks. It provides zero-copy, one-sided, and asynchronous abstractions for efficient communication both within and across nodes.
While the MSCCL++ Primitive API offers fine-grained optimization for GPU experts, MSCCL++ also provides higher-level interfaces that are easier to program. Inspired by MSCCLang (Cowan et al., 2023), we separate the declaration of communication algorithms, done using the MSCCL++ DSL API, from the underlying implementation of the algorithm, done by the MSCCL++ DSL Executor using MSCCL++ primitives. The MSCCL++ DSL accelerates the development of application-specific communication algorithms in new hardware environments. Lastly, for users with the least hardware expertise, we re-implement the popular NCCL API as-is (MSCCL++ Collective API) using MSCCL++ primitives so that applications can adopt it without changing the code at all.
In our evaluation, we compare MSCCL++ with NCCL, ROCm Collective Communication Library (RCCL) (AMD, 2025) and Microsoft Collective Communication Library (MSCCL) (Cowan et al., 2022) on NVIDIA A100, NVIDIA H100, and AMD MI300x GPUs. We show that collective communication kernels like AllReduce and AllGather written using the MSCCL++ DSL API are up to $5.4\times$ faster than the other libraries. We also re-implement some of these communication algorithms directly using the MSCCL++ Primitive API and show that those kernels are on average $3\%$ faster than their counterparts. These algorithm implementations are included in the MSCCL++ Collective library. These benefits translate to AI inference workloads, yielding up to $15\%$ speedup in decodes over NCCL for a popular LLM, Llama2-70b (Touvron et al., 2023).
2. Background and Motivation
2.1. Collective Communication for AI
Modern AI applications are usually based on large machine learning (ML) models that require training and serving on multiple accelerator devices (GPUs in this paper) in a collective manner. They distribute the model and data across multiple GPUs to reduce the computation and memory requirement per GPU. This distribution requires GPUs to share intermediate results at some points during the computation, which is done using collective communication operations. AllReduce is a common collective communication operation that sums the partial results of all GPUs and broadcasts the computed result to all GPUs. AllReduce can be divided into two other collective operations: ReduceScatter and AllGather. ReduceScatter sums the input buffers on all GPUs and distributes the output buffer equally on all GPUs. AllGather collects a distributed buffer from all GPUs and stores the full buffer on each GPU. ReduceScatter and AllGather are often used separately depending on the ML model architectures.
2.2. Existing Hardware Abstractions
Existing collective communication libraries implement conventional networking abstractions such as send and recv inside GPU kernels. This is a straightforward approach to generalize arbitrary communication algorithms in a hardware-agnostic manner. However, it is not the best choice for performance because such software abstractions are not well aligned with the GPU hardware abstractions. Instead, this paper proposes an API design from the perspective of GPU programming, which makes it easier to efficiently utilize GPU hardware. In this section, we provide an overview of the NCCL architecture and discuss its inefficiencies. This is representative of existing collective communication libraries for GPUs, because other popular libraries, such as RCCL (AMD, 2025) and MSCCL (Cowan et al., 2022), are designed based on NCCL and share the same limitations.
2.2.1. NCCL Architecture
NCCL provides a C++ API to setup connections between multiple GPUs and to run communication operations over GPUs. All participating processes (that usually hold one GPU each) of a distributed ML application initialize a NCCL handle to build connections beforehand, and then use the handle to call NCCL kernels (which perform collective operations on GPUs) during the execution. ML frameworks, such as PyTorch (Paszke et al., 2019) and TensorFlow (Abadi et al., 2016), wrap the NCCL API in Python, which are called by an ML model.
1global ringRS(in, nelem, ring)
2 send = ring.sendbuff; recv = ring.recvbuff;
3 sz = ring.buffSz;
4 prim = Primitives<half>
5 (tid, ring.buffSz, ring.prev, ring.next);
6
7 for (off = 0; off < nelem; off += sz)
8 //step 0: push data to the next GPU
9 idx = off + ring.ranks[ring.ranks-1]*sz;
10 prims.copy(in+idx, send, sz);
11 prims.send();
12
13 //k-2 steps: reduce and copy to next GPU
14 for (j = 2; j < ring.ranks; ++j)
15 rankDest = ring.ranks[ring.ranks-j];
16 idx = off + rankDest * buffSz;
17 prims.recv();
18 prim.reduce(in+idx, recv, send, sz, "+");
19 prims.send();
20
21 //step k-1: write the result for this rank
22 idx = off + ring.rank * buffSz;
23 prims.recv();
24 prims.reduce(recv, in+idx, sz, "+");
Figure 1. Ring ReduceScatter (simplified) in NCCL.
NCCL Initialization. During initialization, NCCL obtains following information: (i) the rank of each process, (ii) number of processes, (iii) number of distributed nodes in the cluster, (iv) number of GPUs per node, (v) all links between GPUs in a node, including PCIe, NVLink, and xGMI, and (vi) all links between nodes, like IB and Ethernet. NCCL uses this information to create communication topologies such as ring and tree with bi-directional links. Then, NCCL allocates a send buffer that stores data to send to destination GPU and a receive buffer that stores data received from a source GPU. Instead, users can also register their own send and receive buffers (using ncclMemAlloc()). The metadata of the send and receive buffers on a GPU are shared with all peer GPUs, so that GPUs can either peer-to-peer access buffers of another GPU (through PCIe, NVLink, or xGMI) or use external hardware to copy data (through IB or Ethernet).
NCCL Kernel. NCCL implements a GPU kernel for all collective operations and communication algorithms (ring, tree, etc.). It chooses the best performing algorithm based on the data size of collective communication, namely the message size. NCCL kernels are built using four building block operations, called primitives, which are send, recv, copy, and reduce. (i) send copies data from the send buffer of source GPU to the receive buffer of destination GPU, (ii) recv waits until the transfer by source GPU’s send has finished execution and unblocks send of the source GPU, (iii) copy copies source to destination buffer on the same GPU using GPU threads, and (iv) reduce does elementwise operation, like addition, on two buffers on the same GPU and write the result to another buffer. Calls to send blocks until destination GPUs have called recv to ensure that it is safe to write to send buffer. Additionally, NCCL defines extra operations that execute multiple primitives in a single fused function. However, this does not improve over the limitations we discuss in Section 2.2.2.
As an example, Figure 1 describes the NCCL kernel of a ring-based ReduceScatter for half precision floating point (FP16) using the NCCL primitives. The kernel takes the input buffer, number of elements, a ring of GPUs, and writes output to the input buffer. The kernel obtains pointers to send and receive buffers, size of the buffer, and initializes a primitive object with information about both previous and next GPUs in the ring (lines 25). The main loop of the kernel performs ReduceScatter in a batch of send/receive buffer sizes (lines 724). The loop first copies each batch of input data and sends it to the next GPU (lines 911). The loop then goes through all other ranks in the ring and for each rank (i) computes the corresponding offset in the input buffer for the rank, (ii) receives the partially reduced data from the previous GPU in the receive buffer, (iii) reduces this received data with the input offset and stores it in the send buffer, and (iv) sends this data to the next GPU in the rank (lines 1419). Finally, the GPU stores the data for the offset on this GPU (line 24).
2.2.2. Limitations of NCCL Primitives
The send and recv primitives are not flexible enough to utilize a single-instruction multiple-thread (SIMT) architecture like GPUs efficiently. As GPU threads are executed in groups (warps in CUDA), [^3] Unlike CPU where each thread runs on a separate context and can be context-switched, GPU threads are grouped and executed altogether for a single instruction. Otherwise, it may underutilize the highly-parallel cores and the large memory bandwidth. GPU kernels usually optimize the workload case-by-case to carefully distribute tasks among threads and minimize synchronization overhead. Unfortunately, NCCL only implements a static abstraction that groups 128 - 640 threads to collectively call a single primitive and synchronize them at the end of the call. This static abstraction is not flexible enough to optimize the workload for different links and data sizes, and may introduce unnecessary overheads. We discuss the limitations of NCCL primitives below.
Wasted GPU cycles. The send and recv primitives block the calling threads until the data transfer is completed, and this may waste a bunch of GPU threads’ cycles in a busy-wait while loop. If the underlying link supports peer-to-peer memory access, send directly copies local data into the receive buffer of the destination GPU, which can efficiently utilize parallel GPU threads. However, if the link uses external hardware (e.g., NIC) for the transfer, send only wakes up a CPU thread (writes a few-byte flag) that initiates the transfer and busy-waits until the transfer is completed (see Figure 2). This is a light-weight task that only needs a single GPU thread, and it is a big waste of cycles to block many threads for this. An alternative would be making the send and recv asynchronous, which is also against the existing abstractions in NCCL.
Inflexible Synchronization. To ensure data consistency, the NCCL primitives are designed to be self-synchronous. Such synchronizations are often too conservative to realize correct communication semantics. Consider a loop where $N$ GPUs produce their own data, and each GPU consumes data from all other GPUs, and starts all over again. When a single buffer is written by producers and read by consumers, two types of cross-GPU barriers are required: after each producer to ensure the following remote consumer sees consistent data, and after each consumer to prevent the following remote producer from overwriting what is being read by this consumer. To reduce synchronization overhead, we can use two buffers in rotation and only keep the first type of barriers which is sufficient to prevent overwriting. However, this is not possible with self-synchronized NCCL primitives.
Figure 2. Workflow of the NCCL send and recv primitives between two GPUs connected using InfiniBand (IB). Even though IB NICs can directly read and write GPU memory, we need the CPU to initiate the transfer (ibv_post_send) and wait for the completion (ibv_poll_cq). GPU 0 waits before calling send for next batch of data, and GPU 1 waits before the receive buffer is ready.
Interconnect Optimizations. A link can allow different ways of transferring data. However, NCCL supports only one mode of data transfer for each link. For example, intra-node links, like PCIe and NVLink, have two modes of data transfer: (i) thread-copy, where multiple GPU threads read from source GPU memory and write to destination GPU memory, and (ii) DMA-copy, where the CPU initiates the DMA engine of the GPU to copy from source memory to destination memory using cudaMemcpyDeviceToDevice. Compared to DMA-copy, thread-copy achieves lower latency, but often utilizes less bandwidth. For example, in our AllGather experiments over 8 NVIDIA A100 80G GPUs, thread-copy achieves only up to 227 GB/s NVLink bandwidth, while DMA-copy achieves 263 GB/s (+15.8%). However, NCCL uses only thread-copy unless it is infeasible by hardware. [^4] NCCL uses DMA copy only when peer-to-peer access between GPUs is not supported by the hardware configuration. Moreover, using DMA-copy frees GPU threads to do other work.
2.3. Existing Communication Libraries
NCCL API is designed to provide end-to-end implementations of collective communication operations. This would be what most applications expect from the library, but this is fundamentally limited to support diverse performance-sensitive applications. Like other GPU workloads do, collective communication over GPUs should be differently optimized for different scenarios, depending on the input size, GPU architecture, networking topology, and the application characteristics (Shah et al., 2023; Hwang et al., 2023a). It is very difficult for the library developers to provide the best performance for all cases in a timely manner.
MSCCL (Cowan et al., 2022) tackles this issue by allowing developers to implement their own custom communication algorithms using a DSL (Cowan et al., 2023). However, MSCCL is still based on the NCCL primitives and does not provide enough flexibility to fully utilize the hardware resources, as described in Section 2.2.2.
3. MSCCL++ Overview
Figure 3. Overview of the MSCCL++ stack.
3.1. Hierarchical User Interfaces
MSCCL++ provides three user APIs at hierarchical levels of communication abstractions: Primitive, DSL, and Collective APIs (see Figure 3). The three levels provide different trade-offs in terms of programming effort or expertise and execution latency or performance. The closer the interface is to the hardware, the higher the programmer’s control on execution, so higher the performance and programming effort.
MSCCL++ Primitive API. This replaces NCCL’s primitives with a set of core communication functions that are called from the GPU kernels. This is paired with the bootstrapping API called from the host (CPU) side, which is used for connection setup between GPUs. The primitive interface is the base layer that is used to implement the other two higher-level interfaces. Programmers can also use the primitive interface directly to implement application-specific or hardware-specific optimizations in their own GPU kernels, including those kernels that also perform computation. As the primitive interface is our key contribution, we focus on its details in Section 3 and 4.
MSCCL++ DSL API. We reimplement the MSCCLang (Cowan et al., 2023) DSL over MSCCL++, which allows us to write custom collective communication algorithms in a high-level language. This interface is aimed at users who want to specify communication algorithms or schedules and generate kernels that are optimized for their own workloads (application and hardware). The generated kernel uses the Primitive API by running the DSL Executor, so it may have small runtime overheads.
MSCCL++ Collective API. We reimplement the NCCL API, including its bootstrapping API, as-is over MSCCL++. It is aimed at users with the least expertise; they can simply replace NCCL/RCCL with the MSCCL++ Collective library without changing their application code. This library has the same limitations as that of NCCL (Section 2.3); i.e., it may not provide the best algorithm for certain workloads. For AI inference workloads, we implement the best algorithms in our collective kernels using the MSCCL++ Primitive API.
3.2. Primitive Communication Abstractions
3.2.1. Communication Channels
The supported mode(s) of data transfer differs between interconnects. Therefore, MSCCL++ defines an abstract channel for each data transfer mode supported by the hardware, and exposes communication primitives specific to that channel that can be invoked directly inside a GPU kernel. MSCCL++ defines three communication channels (for readers who are familiar with NCCL, MSCCL++ channel does not correspond to that in NCCL): PortChannel, MemoryChannel, and SwitchChannel, which correspond to port-mapped I/O, memory-mapped I/O, and switch-mapped I/O respectively. A PortChannel uses interconnect ports to communicate between GPUs; i.e., a GPU can initiate the data transfer through a port to another GPU. Ports are controlled by dedicated hardware for I/O, such as DMA engines on GPUs or RDMA NICs. A MemoryChannel uses peer-to-peer memory access to communicate between GPUs; i.e., a GPU can directly access another GPU’s memory. A SwitchChannel uses interconnection-switch-enabled multimem memory access to communicate between GPUs; i.e., a GPU can access the memory of multiple GPUs simultaneously through a switch.
The same interconnect can support multiple modes of data transfer, so multiple channels can be supported on the same interconnect; e.g., all three channels are supported by NVLink, only PortChannel and MemoryChannel are supported by xGMI and PCIe, while only PortChannel is supported by InfiniBand. The implementation of each channel is specific to the interconnect. The SwitchChannel uses NVSwitch multimem instructions that support computation-on-switch. The MemoryChannel uses thread-copy data transfer mode over NVLink, xGMI, and PCIe. On the other hand, the PortChannel uses the DMA-copy data transfer mode over NVLink, xGMI, PCIe, and Infiniband. Current interconnects require a CPU thread to initiate the data transfer; e.g., transferring data over InfiniBand requires the CPU to call ibv_post_send which initiates the RDMA transfer from the sending GPU memory to the receiving GPU memory. Such details are handled transparently by the implementation.
Generality. Port-mapped I/O, memory-mapped I/O, and switch-mapped I/O are complementary I/O methods in general computer architectures. PortChannel, MemoryChannel, and SwitchChannel are our abstractions to provide these I/O methods from inside GPU kernels. We believe that they are flexible enough to accommodate future hardware advances. For example, if future hardware interconnects supports initiating DMA-copy from within a GPU kernel, the same PortChannel API can target the new hardware.
3.2.2. Communication Primitives
All MSCCL++ primitives are defined as a method of a channel. The primitives are all conceptual and do not imply any specific implementation. In this section, we describe the primitives API for the PortChannel with an example. Primitives for the other channels, and the implementation for all the channels are explained in Section 4.
Borrowing the term from MPI (Snir et al., 1998), writing data to the peer’s side through a channel is called put. To provide our communication abstraction as close as possible to hardware capabilities, we design put to be zero-copy (i.e., no intermediate buffers), one-sided (i.e., initiated by a peer without participation of the other peer), and asynchronous. MSCCL++ provides other interfaces for synchronization purposes. Specifically, signal and wait provide a mechanism for synchronization (including memory consistency) across GPUs. flush is for local synchronization that ensures all previous put operations are already on the fly.
Figure 4 shows the semantics of four MSCCL++ primitives: put, signal, wait, and flush. The put primitive transfers data from src0 buffer of GPU-0 to dst1 buffer of GPU-1. GPU-0 calls a following signal primitive that signals GPU-1, which is asynchronous yet strictly ordered with the previous put. GPU-1 has to call the wait primitive before it can read dst1. GPU-0 calls the flush primitive to ensure that the previous put is on the fly and src0 can be safely reused.
MSCCL++ also supports a few fused primitives that can reduce the overhead of API calls. For example, since a put is usually followed by a signal, we provide the putWithSignal function that conducts both at once.
1//async
2put(src0, dst1, size)
3//unsafe to reuse src0
4signal() //async
5flush() //sync
6//safe to reuse src0
(a) GPU-0
//unsafe to read dst1
wait() //sync
//safe to read dst1
(b) GPU-1
Figure 4. MSCCL++ data transfer abstractions. put asynchronously transfers data from one GPU to another. signal and wait synchronize data transfer between GPUs. flush ensures the completion of preceded data transfer.
Example. Figure 5 shows MSCCL++ implementation of the all-pairs ReduceScatter algorithm, which is not used in NCCL. In this algorithm, all GPUs send their data at $\frac{i}{N}$ offset to GPU $i$, the $i^{th}$ GPU does elementwise reduction of the received data, and stores the reduced data. To store the received data, the implementation allocates a scratch buffer per GPU. The kernel’s main loop performs reduction in batches of the scratch size (line 518). On each GPU, the loop first put $\frac{1}{N}$ of data to all other GPUs’ scratch buffer and then signals the completion of transfer (line 910). Then a GPU waits for the data to arrive from other GPUs and reduce it with its part of input data (line 1315). Finally, a barrier among all GPUs is required to ensure scratch is not overwritten for next main loop offset (line 18). Section 5 shows that this algorithm works better than the ring algorithm for smaller data sizes while the ring algorithm works better for large data sizes.
3.2.3. Advantages of Communication Primitives
1global allPairsRS(count, gpus, channels[gpus])
2 sz = channel.scratchSz/gpus.num
3 count = count/gpus.num
4
5 for (off = 0; off < count; off += sz)
6 //Send 1/Nth data to each GPU
7 for (g = 0; g < gpus.num; g++)
8 idx = off + g * count;
9 channels[g].put(idx, sz*g, sz)
10 channels[g].signal()
11
12 //Reduce every pair of GPU
13 for (g = 0; g < gpus.num-1; g++)
14 channels[g].wait()
15 reduce(in + off, channels[g].scratch)
16
17 //barrier on all gpus
18 multiDeviceBarrier();
Figure 5. All-pairs ReduceScatter kernel in MSCCL++. Channels are initialized with source as input and destination as scratch buffer.
The MSCCL++ Primitive API itself does not introduce any novel optimization techniques – instead, it offers new straightforward abstractions that enable quick optimization of GPU communication. The key advantages come from exposing the primitive functionalities as a user interface. First, it allows application developers to easily and quickly optimize GPU communication algorithms for their own infrastructure and workloads. This is valuable in practice because the library developers may not be aware of the specific workloads and hardware configurations of their users. Second, the library developers can quickly deliver new hardware support by implementing the same primitive API for the new hardware. Third, since the primitive API is close to the hardware, it is more flexible to extend or adopt it for new hardware features.
MSCCL++ carefully optimizes GPU communication in the DSL Executor and collective kernels leveraging some of the below features in the Primitive API. Users of the Primitive API can also implement these optimizations directly.
Asynchronous Communication. MSCCL++ contains separate primitives for data transfer and synchronization, allowing asynchronous communication and batching synchronization of multiple data transfers. This abstraction has multiple benefits: (i) allows more algorithms to be specified than the synchronous communication in NCCL, (ii) enables faster collective communication performance by enabling algorithms to combine several synchronizations into one, (iii) frees up resources of peer GPUs for other meaningful work of the application, and (iv) enables better inter-kernel optimizations.
Specialized GPU Kernels. By using channels for specific GPU interconnects, MSCCL++ allows users to specialize kernels for particular interconnects. Furthermore, MSCCL++ avoid extra copies in kernels, leading to a kernel with less code paths and zero register spills. For example, for eight A100 GPUs, the ring AllReduce of NCCL and MSCCL use 94 and 96 registers per thread, respectively, while an MSCCL++ version uses only 32 per thread. Due to these reasons, MSCCL++ kernels execute less instructions, avoid memory accesses due to register spills, and have better instruction cache hits.
Interconnect Optimizations. NCCL only allows a single mode for data transfer over NVLink at a time. In contrast, by using different channel types in MSCCL++ (PortChannel or MemoryChannel), user can utilize both thread-copy and DMA-copy data transfer mode over NVLink.
1class Channel
2 protected:
3 //Src/Dst Buffers set during initialization
4 void* src, *dst;
5 //Semaphore alloced on this GPU
6 uint *semaphore; uint expectedVal;
7 public:
8 //Primitives
9 void signal() ; void wait()
10
11class PortChannel : public Channel
12 public:
13 void put(ulong dstOff, ulong srcOff, ulong sz,
14 uint tid, uint tids)
15
16template<Protocol protocol>
17class MemoryChannel : public Channel
18 public:
19 void put(ulong dstOff, ulong srcOff, ulong sz,
20 uint tid, uint tids, uint flag=0)
21 template<typename T>
22 T read(ulong off, uint tid, uint flag=0)
23
24 template<typename T>
25 void write(ulong off, uint tid, T elem,
26 uint flag=0)
27
28class SwitchChannel : public Channel
29 public:
30 void reduce(ulong dstOff, ulong srcOff,
31 ulong sz, uint tid, uint tids)
32 void broadcast(ulong dstOff, ulong srcOff,
33 ulong sz, uint tid, uint tids)
Figure 6. PortChannel, MemoryChannel, and SwitchChannel primitives in MSCCL++. Implementation of MemoryChannel primitives differ for the type of protocol: LL or HB.
4. MSCCL++ Design and Implementation
4.1. Initialization of MSCCL++
In general, a MSCCL++ program runs as a multi-process on all nodes and GPUs within each node, so that each process runs one GPU. MSCCL++ provides host (CPU) side interfaces for initializing multiple processes, and the interfaces can be specialized by users for their application. The initialization involves setting up connections between GPUs, finding the topology of GPU connections, and constructing channels and buffers.
The first step in initialization is to exchange metadata between host processes, which we call a Bootstrap. It consists of four virtual methods: (i) send to send metadata to another process, (ii) recv to receive metadata from a process, (iii) allGather to collect metadata of all processes, and (iv) barrier to synchronize all processes. MSCCL++ implements a default bootstrap methods using POSIX sockets, while it is straightforward to override with other popular implementations such as MPI or torch.distributed.
The second step is to construct a Communicator that registers communication buffers (either user buffers or dedicated ones) and creates corresponding MSCCL++ channels between GPUs based on the underlying physical links.
4.2. MSCCL++ Primitive API and Channels
A MSCCL++ channel is a connection between two or more GPUs. PortChannel and MemoryChannel are peer-to-peer (i.e., between two GPUs), whereas SwitchChannel is for a group of (i.e., two or more) GPUs. Figure 6 shows three types of channels and its interface in MSCCL++. A channel is initialized with the source and destination buffers, and an integer allocated on the GPU memory that servers as a semaphore.
4.2.1. PortChannel
A PortChannel implements primitives when data transfer is done over ports connected to GPU memory, such as cudaMemcpy for intra-node DMA or ibv_post_send for RDMA. Since data transfer over a port currently requires the CPU to initiate the transfer, each channel creates its own CPU thread. The CPU thread reads data transfer or synchronization requests from a first-in-first-out request queue. The storage, head, and tail of a request queue are allocated using cudaMallocManaged as both CPU and GPU can access it.
We now discuss the workflow of PortChannels using Figure 7. 0 When the GPU calls a primitive, such as put, the first participating thread of the GPU pushes the request to the queue by writing at the head. Before writing to the queue, the GPU checks if the queue is filled, i.e., if the head value is more than the tail value. If the queue is filled then the GPU waits for the CPU to process atleast one request. 1 Then the GPU increments the head to the next element. 2 The CPU thread continuously reads the element at the tail to see if there is a request from the GPU, and when there is a request, the CPU reads the request, zeros out the current element, and increments the tail. 3 Now the CPU thread will process the request. Below we explain the requests generated by primitives and how the CPU handle these requests.
Figure 7. PortChannel workflow for IB from ⓪ when GPU 0 calls put primitive to ⑧ when GPU 1 receives the data.
Data Transfer. put pushes a put request in the queue. 4 The CPU processes this request by starting an RDMA transfer using ibv_post_send. Since this function is asynchronous, the CPU thread immediately returns. When the transfer is happening, peer-GPUs are free to execute code, thereby, improving the overall power efficiency of the system.
Synchronization. 5 signal pushes a signal request in the queue. 6 7 The CPU processes this request by atomically incrementing the semaphore on the receiving GPU using functions, like ibv_atomic_add for IB. The wait primitive on the receiving GPU do not create a request for its CPU, therefore, the receiving CPU is idle, rather wait continuously looks for the semaphore to reach a expected value in a while loop. 8 wait returns after the semaphore is incremented.
Flush. Primitive flush pushes a flush request in the queue and the first thread of the GPU waits until the queue head is equal to or more than the queue tail. The CPU processes the flush request by waiting until all the previous data transfer and synchronization requests have been completed. For example, for IB we use ibv_poll_cq to get the status of all requests. After the flush request is complete, the GPU is free to re-write to the source buffer.
4.2.2. MemoryChannel
A MemoryChannel wraps data transfer methods that use thread-copy mechanism, i.e., directly use GPU threads for writing to peer GPUs memory.
Protocols. The channel provides two protocols, which define data transfer and synchronization technique to tune between low-latency and high-bandwidth: HB protocol provides a high-bandwidth but high-latency protocol, thus, is suitable for larger sizes, and LL protocol provides low-latency but low-bandwidth, thus, is suitable for smaller sizes. Both protocols achieves these properties by synchronizing on the suitable granularity of data.
(i) In the HB protocol, peer GPUs transfer a large chunk of data and synchronize this chunk once using wait and signal primitives. Hence, the synchronization time is amortized over the transfer time leading to high bandwidth. However, since the receiving GPU needs to wait for the whole chunk before processing it, the HB protocol also has high latency.
(ii) In LL protocol, the receiving GPU synchronizes on fewer chunks of data being transferred and process the chunk as soon as it is received, thus, achieving lower latency. However, since the number of synchronization is proportional to the number of elements, this protocol provides lower-bandwidth. The LL protocol works as follows. The put primitive requires an extra integer flag value. For every N1N-1 elements written to the receiving GPU, put also writes the flag. The receiving GPU uses the read primitive, waits until the flag value at $N$ index of receiving buffer is set, and then reads and returns the N1N-1 elements. However, we cannot arbitrarily use any $N$ because GPUs follow a weak memory consistency model where writes to different memory locations by multiple threads can be performed in any order. Therefore, we restrict $N$ to number of elements written by a single instruction, i.e., 4, 8, and 16 bytes memory accesses instructions. The flag value is decided based on the algorithm, such that, all flag values are distinct.
Implementation. A MemoryChannel is initialized with a protocol. A protocol defines the implementation of data transfer and synchronization primitives. We now describe the implementation of primitives based on the protocol as follows:
Data Transfer. Primitive put reads data elements from source buffer and writes to destination buffer starting from different offsets. Similarly, the read primitive reads and returns data from the destination buffer, and the write primitive writes data including the flag for the LL protocol to the destination buffer. To maximize bandwidth for the HB protocol, both primitives use 16-bytes loads and stores. For the LL protocol, both primitives also takes a flag and uses 8-byte load and stores by default or user supplied vector length. Since both put primitives are called by multiple threads (all or first few threads of the kernel), we achieve maximum bandwidth.
Synchronization. Synchronization primitives wait on a semaphore that is an integer allocated on the receiving GPU. The signal primitive atomically increments the receiving GPU’s semaphore and calls threadfence_system to ensure that writes by put and semaphore increments are made available in this order. The wait primitive performs a busy-wait while-loop that checks until the value of semaphore has reached the expected value. This wait is performed by the first thread of the kernel and all other threads wait on a kernel barrier. The channel tracks the expected value using its expectedValue member. The flush primitive is a no-op because after put returns, the source buffer can be reused, even though the write is still in progress.
4.2.3. SwitchChannel
A SwitchChannel provides primitives for performing collective operations among GPUs. These operations usually require specialized hardware support. For example, NVIDIA NVLink 4.0 connects all H100 GPUs to a single NVSwitch and this NVSwitch can perform collective operations including reduce and broadcast. A SwitchChannel in MSCCL++ provides two primitives: reduce to add the corresponding elements of buffers residing on different GPUs and broadcast to send elements from a buffer on a GPU to all other GPUs. We now discuss the implementation of these primitives for NVSwitch, which provides in-network, switch-based aggregation and multicast capabilities using NVLink SHARP (NVLS) technology (NVIDIA, 2025c). However, we believe that these implementations can be generalized to other hardware in the future.
Reduce. This primitive takes a destination buffer as a local address on the local GPU and a source buffer allocated as a multimem address. A multimem address is a virtual address that points to different virtual addresses on each GPU that is a part of the channel/collective. The primitive goes through each element of the destination buffer on the local GPU and executes multimem.ld_reduce PTX instruction using the source element’s multimem address. The multimem instruction fetches the values from all the virtual memory addresses pointed by the multimem address to the switch, does the reduction on the switch, and returns the reduced value to the local GPU. The reduced output is obtained in a register and then written to the destination buffer.
Broadcast. This primitive takes a source buffer as a local address on the local GPU and a destination buffer as a multimem address. The primitive goes through each element of the source buffer on the local GPU, reads the element into a register, and executes the multimem.st PTX instruction using the source register and the destination element’s multimem address. The multimem instruction sends the register value to the switch, which broadcasts and stores the value to all the virtual memory addresses pointed by the multimem address.
4.3. MSCCL++ DSL API and Executor
MSCCL++ DSL API is a Python-based language that allows describing a communication algorithm at a high level. It converts the algorithm description into a sequence of instructions that can be executed by the DSL Executor, which is a GPU kernel that reads and runs instructions back-to-back. MSCCL++ DSL is an extension of the MSCCLang (Cowan et al., 2023) that supports new instructions based on the MSCCL++ Primitive API and lifts a few restrictions of the language; e.g., we enable a single GPU thread block to access multiple GPUs at the same time.
4.4. MSCCL++ Collective API and Kernels
MSCCL++ Collective API is the simplest API that MSCCL++ offers, and it is the same as the popular NCCL API. We implement this API using kernels written in the MSCCL++ Primitive API. Users can either use these default kernels or plug-in their own algorithms written using the MSCCL++ DSL or Primitive APIs. We now briefly describe a few AllReduce algorithms included in the default kernels to: (i) provide ideas on low-level considerations for implementing high-performance GPU communication kernels, and (ii) to show how low-level interfaces can be useful for customization on different workloads.
Env. Name GPU Intra-node Link Network
A100-40G NVIDIA A100 (40G) (8x/node) NVLink 3.0 Mellanox HDR InfiniBand (200 Gb/s, 1x NIC/GPU)
A100-80G NVIDIA A100 (80G) (8x/node) NVLink 3.0 Mellanox HDR InfiniBand (200 Gb/s, 1x NIC/GPU)
H100 NVIDIA H100 (8x/node) NVLink 4.0 Quantum-2 CX7 InfiniBand (400 Gb/s, 1x NIC/GPU)
MI300x AMD MI300x (8x/node) Infinity Fabric Gen 4 Quantum-2 CX7 InfiniBand (400 Gb/s, 1x NIC/GPU)
Table 1. List of environments used for evaluation.
1. One-phase All-pairs (1PA). In all-pairs algorithms, all GPUs concurrently broadcast their own local data to all other GPUs. By one-phase, the algorithm sends all local data to all other GPUs, so that the reduction is done in a single phase. It is suitable for small message sizes where the synchronization overhead between GPUs is more critical than the redundant reduction and data traffic. As we use the 1PA algorithm only for very small message collectives within a single node, we implement only a single version of 1PA that uses the LL protocol with MemoryChannel. MSCCL++ can implement it much more efficiently than other libraries by removing synchronizations.
2. Two-phase All-pairs (2PA). By two-phase, the algorithm splits the AllReduce into two phases: the first for ReduceScatter (each of $N$ GPUs collects and reduces $1/N$ of the data) and the second for AllGather (each GPU broadcasts the reduced data to all other GPUs) (Rabenseifner, 2004). Two-phase algorithms are more bandwidth-efficient and conduct less reduction than one-phase algorithms. In 2PA, the ReduceScatter and AllGather phases are done in the all-pairs manner each. We use 2PA for single-node collectives and implement multiple variants using PortChannel, MemoryChannel with LL or HB protocol, or SwitchChannel. Using MSCCL++, we can optimize 2PA in various ways that are not possible in existing libraries. For example, for up to a few MBs of messages, we exploit rotating buffers to reduce synchronization at the cost of using more memory space. As another example, we can let a single thread group read data from multiple other GPUs at the same time. This allows efficient data reduction compared with other libraries that read data from different GPUs one-by-one, which synchronizes each reduction step.
3. Two-phase Hierarchical (2PH). Hierarchical algorithms exchange minimal data across nodes and do local collectives in each node to complete the operation. It can be faster than all-pairs by reducing the data traffic crossing the nodes. 2PH is a two-phase algorithm that performs ReduceScatter and AllGather in a hierarchical manner each. We use it for multi-node collectives and implement two variants. The first version is for small messages using the LL protocol. Each node conducts local ReduceScatter that splits the data into the number of GPUs in a node. This requires sending more data across nodes and introduces redundant reduction, but it is faster for small messages by reducing synchronization steps. Cross-node communication is done in all-pairs manner. To utilize inter-/intra-node links at the same time, we pipeline the local collective with cross-node communication to overlap. The second version of 2PH is for large messages using HB protocol. This variant also performs all-pairs cross-node communication and local ReduceScatter in a pipelined manner. Unlike the first variant, to utilize the link bandwidth efficiently, the number of data chunks is the same as the number of GPUs.
4.5. Portability and Implementation Effort
MSCCL++ is designed to be portable. Code using MSCCL++ collectives, algorithms using the MSCCL++ DSL, and kernels using the MSCCL++ primitives are portable as long as the underlying MSCCL++ Primitive API supports that hardware. Our design makes it easy to add support for new hardware. As an example, MSCCL++ initially supported only NVIDIA GPUs; to add support for AMD MI300x GPUs, our implementation took only 7 weeks for one developer: 3 weeks for basic AMD GPU support and 4 weeks to develop new AllReduce algorithms that outperform RCCL/MSCCL for message sizes of 1KB - 1GB. It is especially remarkable that the AMD-specific code in MSCCL++ (excluding makefiles and algorithms) is only less than 10 lines of code, while RCCL, the hard-forked NCCL for AMD GPUs, is substantially diverged from NCCL. MSCCL++ could achieve this because the low-level API of AMD GPUs (i.e., HIP) is almost the same as that of NVIDIA GPUs (i.e., CUDA), and the MSCCL++ Primitive API is only a shallow abstraction on top of the low-level API. This significantly reduces the effort to maintain and develop MSCCL++.
The MSCCL++ design makes it easy to extend the Primitive API to support new hardware features. For example, NVIDIA NVLink 4.0 introduced support for multimem instructions for switch-based aggregation and multicast. Since this feature introduced a very different concept from previous interfaces (i.e., PortChannel or MemoryChannel), we developed the SwitchChannel interface to support this feature. The development took only 8 weeks for two developers, including learning the basic usage of this feature, abstracting the feature as a new type of channel (i.e., SwitchChannel), and finally developing a new AllReduce algorithm using SwitchChannel that outperforms NCCL/MSCCL by more than $2.2\times$ in average for message sizes of 1KB - 1GB.
5. Evaluation
Environments. All presented numbers are collected from one of the environments listed in Table 1. Each node is equipped with 8 GPUs per node (either A100, H100, or MI300X), intra-node links between GPUs (either NVLink or Infinity Fabric (a.k.a. xGMI)), and inter-node IB links (one NIC per GPU). All NICs are connected to a single IB networking switch. NVIDIA GPUs use CUDA 12.4, while AMD GPUs use ROCm 6.2. For brevity, the rest of this paper refers to each environment by the name of GPU.
Baselines. We compare MSCCL++ with the state-of-the-art collective communication libraries, including NCCL 2.26.2 (NVIDIA, 2025a), RCCL 2.20.5 (AMD, 2025), [^5] As RCCL partly adopts MSCCL, to prevent confusion, we refer to MSCCL-disabled RCCL. and MSCCL 2.23 (Microsoft, 2025a). MSCCL borrows the stack implementation of NCCL/RCCL and supports custom algorithms. Therefore, MSCCL performs exactly the same as NCCL/RCCL if they use the same algorithm.
5.1. Collective Communication
To compare the performance of collective communication, we implement the AllReduce algorithms described in Section 4.4 using the MSCCL++ DSL API. We also implement various AllGather algorithms in a similar manner, and we omit the implementation details. We present the best number among all implementations for each message size and environment. All NCCL, RCCL, and MSCCL numbers are fine-tuned for each environment and message size by adjusting their environment variables, such as the number of channels (affects the number of threads), chunk size (affects the size of data to be transferred at once), type of algorithm (such as ring, tree, or NVLS (NVIDIA, 2025c)), the topology (XML file that describes the intra-node link topology of GPUs), etc. For MSCCL, we use the fastest algorithm for each message size (Microsoft, 2025b). We use NCCL’s user buffer registration API (i.e., ncclMemAlloc(NVIDIA, 2025f) and the CUDA/HIP Graph APIs for best performance.
For visibility, we separate the range of message sizes into small (up to 1MB, presented latency) and large (1MB and above, presented algorithm bandwidth (AlgoBW) [^6] Defined as the message size divided by the latency.) in the figures. The small message sizes represent inference scenarios (such as LLM token sampling a.k.a. decode (Pope et al., 2023)), while large message sizes represent both training and inference scenarios (such as gradient accumulation during back-propagation, or LLM prompt processing a.k.a. prefill (Pope et al., 2023)).
Results. Figure 8 compares the AllReduce performance over A100-40G. MSCCL++ outperforms baselines for both small messages (up to 4.2x and 3.1x faster over NCCL and MSCCL, respectively) and large messages (up to 1.8x faster over both NCCL and MSCCL). Similarly, Figure 9 compares the AllGather performance over A100-40G. For small messages, MSCCL++ outperforms NCCL and MSCCL by up to 5.4x and 2.3x, respectively, and for large messages, MSCCL++ outperforms NCCL and MSCCL by up to 1.8x and 1.4x, respectively. For AllGather, we observe a few cases where MSCCL slightly outperforms MSCCL++ by up to 8.0%, which is introduced by a performance bug of the DSL API, not by our algorithm implementation. For AllReduce, MSCCL++ is always the fastest.
Figure 8. AllReduce, A100-40G. MnNg means M nodes and N GPUs. X-axis is message size (at the bottom of the figure).
Gain Breakdown. The performance differences can come from both high-level differences in algorithms (i.e., design of data flow routes and pipelines) and low-level optimizations (i.e., how to transfer data, synchronize, and manage buffers). The former benefit is already delivered by MSCCL that enables the customization of algorithms over the NCCL communication stack. Therefore, the key benefits of MSCCL++ lie in the latter by exposing the primitive functionalities to users. Since MSCCL and MSCCL++ use the same high-level algorithm for all cases in Figure 8 and 9, the performance gap between the two comes purely from the benefits of MSCCL++.
NCCL vs MSCCL. MSCCL is faster than NCCL in most cases, and all the benefit comes from better algorithms customized for our environments. The gain is especially big for small message sizes, where MSCCL uses an all-pairs algorithm (which is not supported by NCCL), while NCCL uses the ring algorithm that is worse in terms of latency. For very large message sizes $\geq$ 256MB in multi-node cases, MSCCL achieves higher bandwidth than NCCL because it uses a hierarchical algorithm (2PH), which is more bandwidth-efficient.
Figure 9. AllGather, A100-40G. MnNg means M nodes and N GPUs. X-axis is message size (at the bottom of the figure).
[…]
DSL vs Primitive. From our 9-month experience using the MSCCL++ DSL API, we find that it reduces development time from weeks to days, compared to using the Primitive API directly. The DSL helps automate initialization, check for mistakes, and provide performance analysis. We have implemented both the DSL and Primitive versions up to 2-node experiments to verify that their performance is similar, while 4-node experiments are done only using the DSL. However, as the DSL introduces higher level instructions than the Primitive API, DSL versions perform 3% worse than the Primitive versions on average, and are up to 18% worse in one corner case. Despite limitations, MSCCL++ DSL is still useful for quick prototyping of collective communication and easy-to-understand algorithm description.
Collective vs Primitive. For AI inference workloads, we implement the best algorithms in our collective kernels using the MSCCL++ Primitive API, so there is no performance difference.
Figure 10. Speedup for decodes with tensor parallelism = 8
5.2. LLM Inference
[…]
5.3. Cross-Hardware Support
To show the effectiveness of MSCCL++ on different hardware platforms, Figure 11 and 12 compare the single-node AllReduce performance on H100 and MI300x.
[…]
Figure 11. AllReduce, H100, single-node (8 GPUs).
Figure 12. AllReduce, MI300x, single-node (8 GPUs).
MI300x. MSCCL++ outperforms RCCL and MSCCL by up to 3.8x and 1.9x for small messages, and by up to 2.2x and 1.6x for large messages, respectively. Our algorithm implementations take special considerations to best utilize the Infinity Fabric that peer-to-peer connects all GPUs in a node, unlike NVLink that connects all GPUs to a switch. This means, for best link utilization, we need to copy data to all peers at the same time as much as possible, unlike for NVIDIA GPUs where we can copy data to each peer one-by-one back-to-back. Using the MSCCL++ DSL, this is as easy as changing the order of two nested for loops, one that walks through data elements and the other that walks through different peers.
6. Related Work
NVIDIA Collective Communications Library (NCCL) (NVIDIA, 2025a) and AMD ROCm Collective Communications Library (AMD, 2025) (RCCL) are vendor supplied libraries for NVIDIA and AMD GPUs respectively. Cowan et al. (Cowan et al., 2022) improved over these libraries by allowing execution of collective algorithms specialized for a size and topology. In this paper, we showed that primitives provided by MSCCL++ performs better than the primitives of these libraries.
The primitive interface of MSCCL++ has similarities with some other works in terms of being a GPU-side API for communication. For example, the NVIDIA OpenSHMEM (NVSHMEM) (NVIDIA, 2025b) is a parallel work that provides primitive functions such as nvshmemx_putmem_warp as well as a few collective functions such as nvshmemx_broadcastmem_warp. However, we could not find any implementation where NVSHMEM outperforms NCCL (or MSCCL++) for collective communication, and quantitative comparison is difficult as NVSHMEM is not open source. ARK (Hwang et al., 2023b) is another work that proposes a GPU-side control plane for communication, but it is implemented as a monolithic end-to-end ML system rather than a standalone communication library.
UCX, an open-source framework developed by Shamis et al. (Shamis et al., 2015) and libraries implementing the MPI standard (Thakur et al., 2005) offer a comprehensive set of APIs for high-performance computing (HPC) environments. In contrast, MSCCL++ is more specific to GPU acceleration.
SCCL (Cai et al., 2021), TACCL (Shah et al., 2023), and TE-CCL (Liu et al., 2024) aim to accelerate GPU collective communication by synthesizing efficient data transfer algorithms for performing collective communication, while still using the vendor-provided primitives underneath. The MSCCL++ primitives can be used with the generated algorithms to capture their benefit.
Finally, LLM inference frameworks like vLLM (Kwon et al., 2023) and TensorRT-LLM (NVIDIA, 2025e) implement custom AllReduce kernels for achieving high performance communication. However, the custom implementations are not general-purpose and often only limited to single-node collective communication. MSCCL++, on the other hand, enables communication across multiple types of interconnects and has equivalent AllReduce performance to vLLM’s custom single-node AllReduce.
7. Conclusion
MSCCL++ is a novel GPU communication stack designed for high-performance AI applications. By exposing the primitive communication functionalities as straightforward user interfaces, MSCCL++ enable fine-grained optimizations for GPU experts, while also providing higher-level interfaces for quick optimizations. Moreover, such a design can reduce the overall development and optimization effort for GPU communication, and accelerates adoption of fast evolving hardware technologies. By implementing collective communication using the proposed MSCCL++ interfaces, we can achieve up to $5.4\times$ speedup for standalone collectives, and up to $15\%$ speedup for end-to-end AI inference.

### v4

@@
-MSCCL++: Rethinking GPU Communication Abstractions for Cutting-Edge AI Applications
+MSCCL++: Rethinking GPU Communication Abstractions for AI Inference
 Abstract.
-[…]
-conference: ; ; Now at Microsoft AI. Now at xAI.
+AI applications increasingly run on fast-evolving, heterogeneous hardware to maximize performance, but general-purpose libraries lag in supporting these features.
+Performance-minded programmers often build custom communication stacks that are fast but error-prone and non-portable.
+This paper introduces MSCCL++, a design methodology for developing high-performance, portable communication kernels.
+It provides (1) a low-level, performance-preserving primitive interface that exposes minimal hardware abstractions while hiding the complexities of synchronization and consistency, (2) a higher-level DSL for application developers to implement workload-specific communication algorithms, and (3) a library of efficient algorithms implementing the standard collective API, enabling adoption by users with minimal expertise.
+Compared to state-of-the-art baselines, MSCCL++ achieves geomean speedups of $1.7\times$ (up to $5.4\times$) for collective communication and $1.2\times$ (up to $1.38\times$) for AI inference workloads.
+MSCCL++ is in production of multiple AI services provided by Microsoft Azure, and has also been adopted by RCCL, the GPU collective communication library maintained by AMD.
+MSCCL++ is open source and available at https://github.com/microsoft/mscclpp.
+Our two years of experience with MSCCL++ suggests that its abstractions are robust, enabling support for new hardware features, such as multimem, within weeks of development.
+GPU Communication; Collective Communication; AI Inference; Multi-layered Programming Abstraction
+journalyear: 2026copyright: ccconference: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; March 22–26, 2026; Pittsburgh, PA, USAbooktitle: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’26), March 22–26, 2026, Pittsburgh, PA, USAdoi: 10.1145/3779212.3790188isbn: 979-8-4007-2359-9/2026/03ccs: Software and its engineering Domain specific languagesccs: Software and its engineering API languagesccs: Software and its engineering Requirements analysisccs: Networks Programming interfacesccs: Software and its engineering Abstraction, modeling and modularityccs: Software and its engineering Software performanceccs: Software and its engineering Software usability
 1. Introduction
-GPU communication has become a key area of optimization for high-performance AI applications.
-For instance, modern Large Language Models (LLMs) (Dubey et al., 2024; Brown et al., 2020) are often designed to run a single inference task on tens of GPUs (spanning a few physical nodes) to distribute computation and minimize end-to-end latency.
-Training tasks often scale much further up to tens of thousands of GPUs (Dubey et al., 2024).
-GPUs are connected and communicate with each other through either intra-node (PCIe, NVLink (NVIDIA, 2025d), xGMI (AMD, 2023), etc.) or inter-node (Ethernet or InfiniBand (IB) (inf, 2023)) links.
-Efficient communication is challenging due to the collective communication patterns of AI workloads that frequently distribute and gather data across all GPUs simultaneously.
-For example, communication patterns such as AllReduce, AllGather, AllToAll, Broadcast, and ReduceScatter (Wikipedia, 2025) are frequently used in modern AI applications to distribute computing workload over many GPUs and collect computed results into a single GPU.
-A substantial portion of the end-to-end latency is often attributed to GPU communication in real workloads, usually in the 10% - 40% range of end-to-end LLM workloads.
-For example, a GPT-3 (Brown et al., 2020) model inference may spend 30% of the end-to-end time in AllReduce communication.
-Many efforts have been made to optimize GPU communication, such as developing efficient communication algorithms (routing paths and transfer scheduling) (Shah et al., 2023; Cao et al., 2024; Liu et al., 2024), overlapping communication with computation (Jangda et al., 2022; Wang et al., 2023), and other lower-level optimizations in the stack (NVIDIA, 2025a; AMD, 2025).
-Despite these efforts, achieving high communication performance is still challenging and time consuming in real applications.
-Specifically, practitioners still need to write custom communication code to achieve the best performance, often from scratch.
-For instance, TensorRT-LLM (NVIDIA, 2025e), a popular LLM framework featured by NVIDIA in recent years, implements custom AllReduce communication methods from scratch.
-This is notable given that there already exists a popular GPU communication library developed by NVIDIA for many years, the NVIDIA Collective Communication Library (NCCL) (NVIDIA, 2025a).
-The custom communication of TensorRT-LLM outperforms NCCL in a wide range of LLM scenarios, especially when the data size is relatively small, while TensorRT-LLM still uses NCCL for larger data sizes.
-This raises a simple question: Why build a new ad hoc stack from scratch, instead of developing on top of an existing stack?
-Surprisingly, such ad hoc software development has already become a widely common practice in cutting-edge AI applications.
-From our empirical studies, we find that the root cause lies in the fast evolving hardware.
-The enormous computing demand by modern AI applications is pushing the industry to aggressively upgrade chips and interconnects, resulting in powerful but immature hardware that is dramatically different from the previous generation.
-Cutting-edge AI applications are usually targeted for deployment over the latest hardware to obtain the best performance and efficiency.
-Thus, it is not surprising for existing general-purpose libraries (such as NCCL) to perform sub-optimally for those applications because it takes time for those libraries to optimize for diverse workload scenarios.
-Consequently, production implementations go with custom quick-and-dirty optimization specific to their own workloads and hardware environments.
-This incurs redundant development efforts across applications and writes a lot of non-portable code, which leads to further development effort when a new hardware becomes available.
-In this paper, we discuss an alternative interface design for AI software libraries to support portability and custom optimizations by reducing redundant efforts without harming the flexibility of customization.
-The key idea is to separate high-level abstractions and optimizations from primitive hardware abstractions based on two observations.
-First, existing hardware abstractions like NCCL primitives try to hide hardware features to make it easier to program, but may hinder optimizations.
-Instead, a low-level, primitive interface that minimally abstracts the hardware would make it easier for: (1) application developers to optimize and fine-tune by enabling precise control of the hardware, and (2) library developers to support new hardware features by only providing a shallow layer of abstraction over it.
-Second, existing collective communication libraries (like NCCL) and DSLs (domain-specific languages; like MSCCLang (Cowan et al., 2023)) provide ease of use, but may lack workload-specific optimizations.
-Such higher level abstractions could provide portability and enable users to trade-off performance for programming effort.
-Based on these insights, we propose MSCCL++ (Microsoft Collective Communication Library ++, pronounced em-sickle plus-plus), a novel GPU communication stack designed for high-performance AI applications.
-Unlike existing libraries, MSCCL++ exposes the most fundamental building blocks of communication as a user interface, namely the primitive interface (MSCCL++ Primitive API).
-This interface is very close to GPU hardware as a shallow layer over low-level GPU instructions, thus enabling quick adoption of latest hardware features with low effort.
-The primitive interface abstracts straightforward concepts of communication such as put, signal, wait, and flush.
-This helps translate low-level hardware behaviors into higher-level insights (e.g., bandwidth and latency) for optimizing communication.
-The primitive interface is designed to be very flexible to enable various optimizations that have been difficult to implement with existing stacks.
-It provides zero-copy, one-sided, and asynchronous abstractions for efficient communication both within and across nodes.
-While the MSCCL++ Primitive API offers fine-grained optimization for GPU experts, MSCCL++ also provides higher-level interfaces that are easier to program.
-Inspired by MSCCLang (Cowan et al., 2023), we separate the declaration of communication algorithms, done using the MSCCL++ DSL API, from the underlying implementation of the algorithm, done by the MSCCL++ DSL Executor using MSCCL++ primitives.
-The MSCCL++ DSL accelerates the development of application-specific communication algorithms in new hardware environments.
-Lastly, for users with the least hardware expertise, we re-implement the popular NCCL API as-is (MSCCL++ Collective API) using MSCCL++ primitives so that applications can adopt it without changing the code at all.
+GPU communication has emerged as a critical bottleneck and an opportunity for optimization in AI systems.
+For instance, communication kernels account for approximately 10%-40% in many real-world Large Language Model (LLM) inference tasks (NVIDIA, 2025f; Gond et al., 2025; Hwang et al., 2023a).
+Recent research has addressed communication bottlenecks, including developing efficient communication algorithms (e.g., routing paths and transfer scheduling) (Shah et al., 2023; Cao et al., 2024; Liu et al., 2024), overlapping communication with computation (Jangda et al., 2022; Wang et al., 2023), and other lower-level optimizations in the stack (NVIDIA, 2025d; AMD, 2025).
+Despite these efforts, achieving high communication performance is still challenging and time-consuming in real applications.
+A common practice today is for practitioners to write custom communication code from scratch to achieve maximum performance without using standard libraries such as the NVIDIA Collective Communication Library (NCCL) (NVIDIA, 2025d).
+For instance, NVIDIA’s TensorRT-LLM (NVIDIA, 2025g) implements custom AllReduce communication kernels rather than relying on or updating NCCL.
+There are multiple reasons for doing so.
+First, achieving peak performance requires a careful balance between latency and bandwidth, tailored to the specific message sizes and communication patterns of a workload — something a general-purpose library, designed to support many workloads, cannot always optimize (Cowan et al., 2023).
+For example, TensorRT-LLM uses a custom AllReduce that outperforms NCCL on small data sizes, while falling back to NCCL for larger ones.
+Second, communication collectives often involve computations (e.g., reductions), and their performance depends on exploiting parallelism in ways that interact well with the necessary computation, further motivating workload-specific customization.
+Third, the enormous computational demand of modern AI applications is driving rapid evolution in both chips and interconnects, and practitioners rush to exploit these new capabilities long before general-purpose libraries such as NCCL are fully adapted.
+The practice of writing custom communication code comes at a huge engineering cost.
+Developing these libraries is challenging and error prone.
+New hardware capabilities are often immature or poorly documented, forcing developers to navigate low-level details with limited guidance.
+There are also complex synchronization and data consistency challenges that arise from interactions across GPUs, CPUs, and NICs, which must be carefully managed to avoid correctness issues and performance bottlenecks.
+Finally, achieving maximum efficiency typically demands the design of custom communication algorithms tailored to the specific workload and hardware, further increasing development complexity.
+Our hypothesis to address these issues is that multi-layered programming abstractions are needed to provide performance, portability, and productivity at the same time.
+This paper proposes MSCCL++ [^1] MSCCL++stands for Microsoft Collective Communication Library ++, pronounced em-sickle plus-plus., a design methodology that evaluates this hypothesis for high-performance GPU communication.
+MSCCL++ provides portable, hierarchical abstractions that enable users to explore the trade-offs between productivity and performance.
+At its core, MSCCL++ provides a Primitive API that exposes minimal performance-preserving abstractions of hardware interconnect methods.
+These abstractions provide full flexibility for customization, including fine-grained overlapping of compute and communication, while hiding the complexities of synchronization and ensuring portability across hardware vendors and generations.
+Building on the Primitive API, MSCCL++ provides a high-level Domain-Specific Language (DSL) for specifying custom communication algorithms while the MSCCL++ DSL Executor handles implementation concerns and performs optimizations when executing the DSL program using the Primitive API.
+Using the DSL, MSCCL++ provides a library of collective algorithms optimized for different scenarios that together provide the same API as NCCL.
+Our design is thus based on a separation of concerns: (1) the Primitive API provides a portable substrate for expert programmers to exploit hardware-specific features for the best performance; (2) the DSL API offers a streamlined interface that allows application developers to implement workload-specific and network-topology-specific optimizations with minimal effort; and (3) the Collective API enables users to leverage these specialized implementations with no additional effort.
+These three layers offer abstractions that improve programmability while enabling progressive customization layer-by-layer to get the last ounce of performance.
+Prior primitive interfaces not only are hardware-vendor-specific but also often mask certain hardware features, limiting opportunities for optimization; e.g., NCCL provides a low-level send-recv abstraction that prohibits developers from exploiting one-sided, asynchronous communication over interconnects, and NVSHMEM (NVIDIA, 2025e) provides a low-level shared-memory abstraction that prevents interconnect-specific optimizations such as using port-mapped DMA-copy on NVLink.
+In contrast, the MSCCL++ Primitive API exposes the fundamental building blocks of communication for different data transfer modes supported by hardware interconnects: memory-mapped I/O, port-mapped I/O, and switch-mapped I/O.
+While the building blocks are specific to the data transfer modes, they are portable across hardware vendors and generations supporting the corresponding data transfer mode.
+This interface supports one-sided and asynchronous abstractions for efficient communication both within and across nodes.
+The implementation of these primitives handles orchestration and synchronization across GPU, CPU, and NIC, providing clear data consistency and synchronization semantics.
+The MSCCL++ DSL allows developers to specify communication algorithms while abstracting away implementation details.
+To improve programmability, it provides a thread-block-based API with a global view across all GPUs and exposes the underlying one-sided, asynchronous communication abstractions of the Primitive API.
+MSCCL++ automatically inserts intra-thread-block synchronizations, fuses instructions, and removes unnecessary memory accesses before lowering to an execution plan that is interpreted by the MSCCL++ DSL Executor.
+This design accelerates the development of application-specific communication algorithms for new hardware environments.
+Unlike MSCCLang (Cowan et al., 2023), our DSL retains the one-sided and asynchronous properties of the Primitive API and enables algorithms that overlap communication with computation.
+Using the DSL, we implement efficient communication algorithms, including those with computation-communication overlap, that are tuned for AI inference workloads.
+These algorithms are bundled into our library that provides the MSCCL++ Collective API.
 In our evaluation, we compare MSCCL++ with NCCL, ROCm Collective Communication Library (RCCL) (AMD, 2025) and Microsoft Collective Communication Library (MSCCL) (Cowan et al., 2022) on NVIDIA A100, NVIDIA H100, and AMD MI300x GPUs.
-We show that collective communication kernels like AllReduce and AllGather written using the MSCCL++ DSL API are up to $5.4\times$ faster than the other libraries.
-We also re-implement some of these communication algorithms directly using the MSCCL++ Primitive API and show that those kernels are on average $3\%$ faster than their counterparts.
-These algorithm implementations are included in the MSCCL++ Collective library.
-These benefits translate to AI inference workloads, yielding up to $15\%$ speedup in decodes over NCCL for a popular LLM, Llama2-70b (Touvron et al., 2023).
+We show that collective communication kernels like AllReduce and AllGather implemented using the MSCCL++ DSL are on average $1.99\times$, $2.08\times$, and $1.43\times$ faster than NCCL, RCCL, and MSCCL respectively.
+We also develop some of these communication algorithms directly using the MSCCL++ Primitive API and show that those kernels are on average $3\%$ faster than their DSL counterparts.
+These benefits translate to AI inference workloads.
+LLM inference decodes are on average $1.11\times$ and $1.31\times$ faster by using MSCCL++ instead of NCCL in vLLM (Kwon et al., 2023) and SGLang (SGLang, 2025) respectively.
+The MSCCL++ code base (Microsoft, 2025c) has been publicly available for two years.
+During this time, RCCL (AMD, 2025) has adopted both the MSCCL++ APIs and library as the default to support collective communication primitives on current and upcoming AMD hardware.
+SGLang (SGLang, 2025), a popular LLM inference framework, uses the MSCCL++ library for its collective communication.
+As a further validation of our design, when NVIDIA introduced support for multimem instructions for switch-mapped I/O (aggregation and multicast), incorporating this support in MSCCL++ (Section 4.3) required only $16$ person-weeks, allowing the rest of the MSCCL++ stack to leverage these primitives.
+Similarly, supporting NVIDIA’s multi-node NVLink (NVIDIA, 2025b) took only $2$ person-weeks.
 2. Background and Motivation
-2.1. Collective Communication for AI
-Modern AI applications are usually based on large machine learning (ML) models that require training and serving on multiple accelerator devices (GPUs in this paper) in a collective manner.
-They distribute the model and data across multiple GPUs to reduce the computation and memory requirement per GPU.
-This distribution requires GPUs to share intermediate results at some points during the computation, which is done using collective communication operations.
+2.1. Collective Communications
+Communication in AI workloads often happens through collective operations in which multiple GPUs coordinate to transfer data among themselves and optionally performing computations on the data.
 AllReduce is a common collective communication operation that sums the partial results of all GPUs and broadcasts the computed result to all GPUs.
@@
 AllGather collects a distributed buffer from all GPUs and stores the full buffer on each GPU.
-ReduceScatter and AllGather are often used separately depending on the ML model architectures.
-2.2. Existing Hardware Abstractions
-Existing collective communication libraries implement conventional networking abstractions such as send and recv inside GPU kernels.
-This is a straightforward approach to generalize arbitrary communication algorithms in a hardware-agnostic manner.
-However, it is not the best choice for performance because such software abstractions are not well aligned with the GPU hardware abstractions.
-Instead, this paper proposes an API design from the perspective of GPU programming, which makes it easier to efficiently utilize GPU hardware.
-In this section, we provide an overview of the NCCL architecture and discuss its inefficiencies.
-This is representative of existing collective communication libraries for GPUs, because other popular libraries, such as RCCL (AMD, 2025) and MSCCL (Cowan et al., 2022), are designed based on NCCL and share the same limitations.
-2.2.1. NCCL Architecture
-NCCL provides a C++ API to setup connections between multiple GPUs and to run communication operations over GPUs.
-All participating processes (that usually hold one GPU each) of a distributed ML application initialize a NCCL handle to build connections beforehand, and then use the handle to call NCCL kernels (which perform collective operations on GPUs) during the execution.
-ML frameworks, such as PyTorch (Paszke et al., 2019) and TensorFlow (Abadi et al., 2016), wrap the NCCL API in Python, which are called by an ML model.
-1global ringRS(in, nelem, ring)
-2 send = ring.sendbuff; recv = ring.recvbuff;
-3 sz = ring.buffSz;
-4 prim = Primitives<half>
-5 (tid, ring.buffSz, ring.prev, ring.next);
-6
-7 for (off = 0; off < nelem; off += sz)
-8 //step 0: push data to the next GPU
-9 idx = off + ring.ranks[ring.ranks-1]*sz;
-10 prims.copy(in+idx, send, sz);
-11 prims.send();
-12
-13 //k-2 steps: reduce and copy to next GPU
-14 for (j = 2; j < ring.ranks; ++j)
-15 rankDest = ring.ranks[ring.ranks-j];
-16 idx = off + rankDest * buffSz;
-17 prims.recv();
-18 prim.reduce(in+idx, recv, send, sz, "+");
-19 prims.send();
-20
-21 //step k-1: write the result for this rank
-22 idx = off + ring.rank * buffSz;
-23 prims.recv();
-24 prims.reduce(recv, in+idx, sz, "+");
-Figure 1. Ring ReduceScatter (simplified) in NCCL.
-NCCL Initialization.
-During initialization, NCCL obtains following information: (i) the rank of each process, (ii) number of processes, (iii) number of distributed nodes in the cluster, (iv) number of GPUs per node, (v) all links between GPUs in a node, including PCIe, NVLink, and xGMI, and (vi) all links between nodes, like IB and Ethernet.
-NCCL uses this information to create communication topologies such as ring and tree with bi-directional links.
-Then, NCCL allocates a send buffer that stores data to send to destination GPU and a receive buffer that stores data received from a source GPU.
-Instead, users can also register their own send and receive buffers (using ncclMemAlloc()).
-The metadata of the send and receive buffers on a GPU are shared with all peer GPUs, so that GPUs can either peer-to-peer access buffers of another GPU (through PCIe, NVLink, or xGMI) or use external hardware to copy data (through IB or Ethernet).
-NCCL Kernel.
-NCCL implements a GPU kernel for all collective operations and communication algorithms (ring, tree, etc.).
-It chooses the best performing algorithm based on the data size of collective communication, namely the message size.
-NCCL kernels are built using four building block operations, called primitives, which are send, recv, copy, and reduce.
-(i) send copies data from the send buffer of source GPU to the receive buffer of destination GPU, (ii) recv waits until the transfer by source GPU’s send has finished execution and unblocks send of the source GPU, (iii) copy copies source to destination buffer on the same GPU using GPU threads, and (iv) reduce does elementwise operation, like addition, on two buffers on the same GPU and write the result to another buffer.
-Calls to send blocks until destination GPUs have called recv to ensure that it is safe to write to send buffer.
-Additionally, NCCL defines extra operations that execute multiple primitives in a single fused function.
-However, this does not improve over the limitations we discuss in Section 2.2.2.
-As an example, Figure 1 describes the NCCL kernel of a ring-based ReduceScatter for half precision floating point (FP16) using the NCCL primitives.
-The kernel takes the input buffer, number of elements, a ring of GPUs, and writes output to the input buffer.
-The kernel obtains pointers to send and receive buffers, size of the buffer, and initializes a primitive object with information about both previous and next GPUs in the ring (lines 25).
-The main loop of the kernel performs ReduceScatter in a batch of send/receive buffer sizes (lines 724).
-The loop first copies each batch of input data and sends it to the next GPU (lines 911).
-The loop then goes through all other ranks in the ring and for each rank (i) computes the corresponding offset in the input buffer for the rank, (ii) receives the partially reduced data from the previous GPU in the receive buffer, (iii) reduces this received data with the input offset and stores it in the send buffer, and (iv) sends this data to the next GPU in the rank (lines 1419).
-Finally, the GPU stores the data for the offset on this GPU (line 24).
-2.2.2. Limitations of NCCL Primitives
-The send and recv primitives are not flexible enough to utilize a single-instruction multiple-thread (SIMT) architecture like GPUs efficiently.
-As GPU threads are executed in groups (warps in CUDA), [^3] Unlike CPU where each thread runs on a separate context and can be context-switched, GPU threads are grouped and executed altogether for a single instruction.
-Otherwise, it may underutilize the highly-parallel cores and the large memory bandwidth.
-GPU kernels usually optimize the workload case-by-case to carefully distribute tasks among threads and minimize synchronization overhead.
-Unfortunately, NCCL only implements a static abstraction that groups 128 - 640 threads to collectively call a single primitive and synchronize them at the end of the call.
-This static abstraction is not flexible enough to optimize the workload for different links and data sizes, and may introduce unnecessary overheads.
-We discuss the limitations of NCCL primitives below.
-Wasted GPU cycles.
-The send and recv primitives block the calling threads until the data transfer is completed, and this may waste a bunch of GPU threads’ cycles in a busy-wait while loop.
-If the underlying link supports peer-to-peer memory access, send directly copies local data into the receive buffer of the destination GPU, which can efficiently utilize parallel GPU threads.
-However, if the link uses external hardware (e.g., NIC) for the transfer, send only wakes up a CPU thread (writes a few-byte flag) that initiates the transfer and busy-waits until the transfer is completed (see Figure 2).
-This is a light-weight task that only needs a single GPU thread, and it is a big waste of cycles to block many threads for this.
-An alternative would be making the send and recv asynchronous, which is also against the existing abstractions in NCCL.
-Inflexible Synchronization.
-To ensure data consistency, the NCCL primitives are designed to be self-synchronous.
-Such synchronizations are often too conservative to realize correct communication semantics.
-Consider a loop where $N$ GPUs produce their own data, and each GPU consumes data from all other GPUs, and starts all over again.
-When a single buffer is written by producers and read by consumers, two types of cross-GPU barriers are required: after each producer to ensure the following remote consumer sees consistent data, and after each consumer to prevent the following remote producer from overwriting what is being read by this consumer.
-To reduce synchronization overhead, we can use two buffers in rotation and only keep the first type of barriers which is sufficient to prevent overwriting.
-However, this is not possible with self-synchronized NCCL primitives.
-Figure 2.
-Workflow of the NCCL send and recv primitives between two GPUs connected using InfiniBand (IB).
-Even though IB NICs can directly read and write GPU memory, we need the CPU to initiate the transfer (ibv_post_send) and wait for the completion (ibv_poll_cq).
-GPU 0 waits before calling send for next batch of data, and GPU 1 waits before the receive buffer is ready.
-Interconnect Optimizations.
-A link can allow different ways of transferring data.
-However, NCCL supports only one mode of data transfer for each link.
-For example, intra-node links, like PCIe and NVLink, have two modes of data transfer: (i) thread-copy, where multiple GPU threads read from source GPU memory and write to destination GPU memory, and (ii) DMA-copy, where the CPU initiates the DMA engine of the GPU to copy from source memory to destination memory using cudaMemcpyDeviceToDevice.
-Compared to DMA-copy, thread-copy achieves lower latency, but often utilizes less bandwidth.
-For example, in our AllGather experiments over 8 NVIDIA A100 80G GPUs, thread-copy achieves only up to 227 GB/s NVLink bandwidth, while DMA-copy achieves 263 GB/s (+15.8%).
-However, NCCL uses only thread-copy unless it is infeasible by hardware.
-[^4] NCCL uses DMA copy only when peer-to-peer access between GPUs is not supported by the hardware configuration.
-Moreover, using DMA-copy frees GPU threads to do other work.
-2.3. Existing Communication Libraries
-NCCL API is designed to provide end-to-end implementations of collective communication operations.
-This would be what most applications expect from the library, but this is fundamentally limited to support diverse performance-sensitive applications.
-Like other GPU workloads do, collective communication over GPUs should be differently optimized for different scenarios, depending on the input size, GPU architecture, networking topology, and the application characteristics (Shah et al., 2023; Hwang et al., 2023a).
-It is very difficult for the library developers to provide the best performance for all cases in a timely manner.
-MSCCL (Cowan et al., 2022) tackles this issue by allowing developers to implement their own custom communication algorithms using a DSL (Cowan et al., 2023).
-However, MSCCL is still based on the NCCL primitives and does not provide enough flexibility to fully utilize the hardware resources, as described in Section 2.2.2.
+2.2. Vendor Communication Libraries
+GPU vendors provide communication libraries that implement the collective operations.
+NVIDIA Collective Communication Libraries (NCCL) (NVIDIA, 2025d) is an optimized library for collective communications on NVIDIA hardware.
+Similarly, AMD ROCm Collective Communication Library (RCCL) (AMD, 2025) is optimized for AMD hardware.
+Microsoft Collective Communication Library (MSCCL) (Cowan et al., 2022) enables executing custom communication algorithm optimized for a particular input size and topology of the hardware.
+We briefly discuss the design of NCCL (see Hu et al.
+(Hu et al., 2025) for a detailed description), since both RCCL and MSCCL are based on NCCL and they both have the same limitations as NCCL.
+[^2] RCCL has since adopted code and abstractions from MSCCL++.
+NCCL implements a GPU kernel for all collective operations.
+The implementation chooses the communication algorithms (ring, tree, etc.) to implement the collective based on the data size.
+The core communication happens through four primitive operations: send, recv, copy, and reduce.
+NCCL allocates send and receive buffers to facilitate the transfer.
+These operations are synchronous.
+For instance, calls to send blocks until destination GPUs have called recv to ensure that it is safe to overwrite the send buffer.
+2.3. Limitations of NCCL Primitives
+NCCL primitives are synchronous, which improves programmability, but wastes GPU cycles in busy-wait loops and prevents overlapping of computation and communication.
+In addition, NCCL’s send and recv primitives work over internal send/receive buffers incurring local memory copies.
+Finally, by being a general-purpose library, NCCL hard codes many performance optimizations providing inflexibility to the users.
+For instance, NCCL provides limited capability to control the parallelism for performing reduction or copy operations during communication.
+Similarly, NCCL hardcodes a single transfer mode per link, even though interconnects like PCIe and NVLink support multiple modes of transfer with subtle performance tradeoffs.
 3. MSCCL++Overview
-Figure 3. Overview of the MSCCL++ stack.
-3.1. Hierarchical User Interfaces
 MSCCL++provides three user APIs at hierarchical levels of communication abstractions: Primitive, DSL, and Collective APIs (see Figure 1).
@@
 The closer the interface is to the hardware, the higher the programmer’s control on execution, so higher the performance and programming effort.
+Figure 1. Overview of the MSCCL++ stack.
 MSCCL++ Primitive API.
@@
 Programmers can also use the primitive interface directly to implement application-specific or hardware-specific optimizations in their own GPU kernels, including those kernels that also perform computation.
-As the primitive interface is our key contribution, we focus on its details in Section 3 and 4.
 MSCCL++ DSL API.
-We reimplement the MSCCLang (Cowan et al., 2023) DSL over MSCCL++, which allows us to write custom collective communication algorithms in a high-level language.
+Based on the primitive interface, we introduce a new DSL to write custom collective communication algorithms in a high-level language.
 This interface is aimed at users who want to specify communication algorithms or schedules and generate kernels that are optimized for their own workloads (application and hardware).
@@
 MSCCL++ Collective API.
-We reimplement the NCCL API, including its bootstrapping API, as-is over MSCCL++.
+We re-implement the NCCL API, including its bootstrapping API, as-is over MSCCL++.
 It is aimed at users with the least expertise; they can simply replace NCCL/RCCL with the MSCCL++ Collective library without changing their application code.
-This library has the same limitations as that of NCCL (Section 2.3); i.e., it may not provide the best algorithm for certain workloads.
-For AI inference workloads, we implement the best algorithms in our collective kernels using the MSCCL++ Primitive API.
-3.2. Primitive Communication Abstractions
-3.2.1. Communication Channels
-The supported mode(s) of data transfer differs between interconnects.
-Therefore, MSCCL++ defines an abstract channel for each data transfer mode supported by the hardware, and exposes communication primitives specific to that channel that can be invoked directly inside a GPU kernel.
-MSCCL++ defines three communication channels (for readers who are familiar with NCCL, MSCCL++ channel does not correspond to that in NCCL): PortChannel, MemoryChannel, and SwitchChannel, which correspond to port-mapped I/O, memory-mapped I/O, and switch-mapped I/O respectively.
-A PortChannel uses interconnect ports to communicate between GPUs; i.e., a GPU can initiate the data transfer through a port to another GPU.
-Ports are controlled by dedicated hardware for I/O, such as DMA engines on GPUs or RDMA NICs.
-A MemoryChannel uses peer-to-peer memory access to communicate between GPUs; i.e., a GPU can directly access another GPU’s memory.
-A SwitchChannel uses interconnection-switch-enabled multimem memory access to communicate between GPUs; i.e., a GPU can access the memory of multiple GPUs simultaneously through a switch.
-The same interconnect can support multiple modes of data transfer, so multiple channels can be supported on the same interconnect; e.g., all three channels are supported by NVLink, only PortChannel and MemoryChannel are supported by xGMI and PCIe, while only PortChannel is supported by InfiniBand.
-The implementation of each channel is specific to the interconnect.
-The SwitchChannel uses NVSwitch multimem instructions that support computation-on-switch.
-The MemoryChannel uses thread-copy data transfer mode over NVLink, xGMI, and PCIe.
-On the other hand, the PortChannel uses the DMA-copy data transfer mode over NVLink, xGMI, PCIe, and Infiniband.
-Current interconnects require a CPU thread to initiate the data transfer; e.g., transferring data over InfiniBand requires the CPU to call ibv_post_send which initiates the RDMA transfer from the sending GPU memory to the receiving GPU memory.
-Such details are handled transparently by the implementation.
-Generality.
-Port-mapped I/O, memory-mapped I/O, and switch-mapped I/O are complementary I/O methods in general computer architectures.
+Similar to NCCL/RCCL, this library may not provide the best algorithm for certain workloads.
+For AI inference workloads, we implement the best algorithms in our collective kernels using the MSCCL++ DSL.
+4. MSCCL++Primitive Abstractions
+1class PortChannel
+2 void *src, *dst; // set during init
+3 uint *semaphore; uint expectedVal;
+4
+5 void put(ulong dstOff, ulong srcOff, ulong sz);
+6 void signal();
+7 void wait();
+8 void flush();
+9
+10template<Protocol protocol> // LL or HB protocol
+11class MemoryChannel
+12 void *src, *dst; // set during init
+13 uint *semaphore; uint expectedVal;
+14
+15 void put(ulong dstOff, ulong srcOff, ulong sz,
+16 uint tid, uint tids, uint flag=0);
+17 template<typename T>
+18 T read(ulong off, uint flag=0);
+19 template<typename T>
+20 void write(ulong off, T elem, uint flag=0);
+21 void signal();
+22 void wait();
+23 void flush();
+24
+25class SwitchChannel
+26 void *local, *multimem; // set during init
+27
+28 void reduce(ulong dstOff,ulong srcOff,ulong sz);
+29 void broadcast(ulong dstOff,ulong srcOff,
+30 ulong sz);
+Figure 2. Channel abstractions in MSCCL++ Primitive API.
+MSCCL++Primitive API provides the minimal hardware abstractions for GPU communication.
+These abstractions are close to the hardware to avoid performance overheads.
+At the same time, they hide the complexities of GPU/CPU/NIC orchestration, synchronization, and memory consistency issues to simplify the programmability.
+The modes of data transfer differ across different interconnects.
+At its core, MSCCL++ provides the notion of a Channel to capture these modes of communication with basic synchronization primitives.
+Specifically, MSCCL++ provides (1) PortChannel for port-mapped I/O allowing GPUs to initiate data transfer through dedicated hardware I/O ports such as DMA engines or RDMA NICs, (2) MemoryChannel for memory-mapped I/O with direct peer-to-peer memory access across GPUs, and (3) SwitchChannel for switch-mapped I/O using multicasting memory access enabled by interconnect switches.
+Each of these channels exposes additional primitives specific to that channel.
+An interconnect might support multiple modes of data transfer and hence can support multiple of these interfaces.
+For instance, NVLink supports all the three, while xGMI and PCIe only support port-mapped and memory-mapped I/O.
+Our intent is that these abstractions are general allowing the library implementer to quickly adapt to changing hardware capabilities.
+For instance, the current implementation of PortChannel for NVLink requires coordination with the CPU to initiate a DMA engine (cudaMemcpy), but can directly initiate it from the GPU if future hardware supports it.
+Port-mapped, memory-mapped, and switch-mapped I/O are complementary I/O methods in general computer architectures.
 PortChannel, MemoryChannel, and SwitchChannel are our abstractions to provide these I/O methods from inside GPU kernels.
 We believe that they are flexible enough to accommodate future hardware advances.
-For example, if future hardware interconnects supports initiating DMA-copy from within a GPU kernel, the same PortChannel API can target the new hardware.
-3.2.2. Communication Primitives
-All MSCCL++ primitives are defined as a method of a channel.
-The primitives are all conceptual and do not imply any specific implementation.
-In this section, we describe the primitives API for the PortChannel with an example.
-Primitives for the other channels, and the implementation for all the channels are explained in Section 4.
-Borrowing the term from MPI (Snir et al., 1998), writing data to the peer’s side through a channel is called put.
+For example, if future hardware interconnects support initiating DMA-copy from within a GPU kernel (Hwang et al., 2023b), the same PortChannel API can target the new hardware.
+Figure 2 shows three types of channels and its interface in MSCCL++.
+MSCCL++ also supports a few fused primitives (omitted in Figure 2 for brevity) that can reduce the overhead of API calls.
+For example, since a put is usually followed by a signal, we provide the put_with_signal function that performs both at once.
+The buffers and semaphores are allocated on the GPU memory before initialization of the channels.
+Next, we describe the three kinds of MSCCL++ channels.
+4.1. PortChannel
+A PortChannel implements primitives when data transfer is done over ports connected to GPU memory, such as cudaMemcpy for intra-node DMA or ibv_post_send for RDMA.
+Borrowing the term from MPI (Snir et al., 1998), writing data to the peer’s side is called put.
 To provide our communication abstraction as close as possible to hardware capabilities, we design put to be zero-copy (i.e., no intermediate buffers), one-sided (i.e., initiated by a peer without participation of the other peer), and asynchronous.
-MSCCL++ provides other interfaces for synchronization purposes.
-Specifically, signal and wait provide a mechanism for synchronization (including memory consistency) across GPUs.
-flush is for local synchronization that ensures all previous put operations are already on the fly.
-Figure 4 shows the semantics of four MSCCL++ primitives: put, signal, wait, and flush.
+Figure 3 shows the semantics of the synchronization primitives: signal, wait, and flush.
 The put primitive transfers data from src0 buffer of GPU-0 to dst1 buffer of GPU-1.
-GPU-0 calls a following signal primitive that signals GPU-1, which is asynchronous yet strictly ordered with the previous put.
-GPU-1 has to call the wait primitive before it can read dst1.
-GPU-0 calls the flush primitive to ensure that the previous put is on the fly and src0 can be safely reused.
-MSCCL++ also supports a few fused primitives that can reduce the overhead of API calls.
-For example, since a put is usually followed by a signal, we provide the putWithSignal function that conducts both at once.
-1//async
-2put(src0, dst1, size)
-3//unsafe to reuse src0
-4signal() //async
-5flush() //sync
-6//safe to reuse src0
+The signal is an asynchronous operation that is strictly ordered with respect to the previous transfer operations, such as put here.
+The matching wait primitive called by GPU-1 blocks till it receives GPU-0’s signal at which time GPU-1 is allowed to access the transferred data in dst1.
+GPU-0 calls the flush primitive to ensure that the transfer is complete from its perspective allowing it to reuse the src0 buffer.
+Implementation.
+Since data transfer over a port currently requires the CPU to initiate the transfer, each channel creates its own CPU thread.
+The CPU thread reads data transfer or synchronization requests from a first-in-first-out request queue.
+The storage and tail of a request queue are allocated using cudaMallocManaged as both CPU and GPU can access it, while the head is on GPU and only accessed by GPU.
+We now discuss the workflow of PortChannels using Figure 4.
+0 When the GPU calls a primitive, such as put, the first participating thread of the GPU pushes the request to the queue by writing at the head.
+Before writing to the queue, the GPU checks if the queue is filled, i.e., if the head value is more than the tail value.
+If the queue is filled then the GPU waits for the CPU to process at least one request.
+1 Then the GPU increments the head to the next element.
+2 The CPU thread continuously reads the element at the tail to see if there is a request from the GPU, and when there is a request, the CPU reads the request, zeros out the current element, and increments the tail.
+3 Now the CPU thread will process the request.
+Below we explain three types of requests and how the CPU handles them over InfiniBand (IB) as an example.
+//async
+put(src0, dst1, size)
+//unsafe to reuse src0
+signal() //async
+flush() //sync
+//safe to reuse src0
 (a) GPU-0
@@
 flush ensures the completion of preceded data transfer.
-Example.
-Figure 5 shows MSCCL++ implementation of the all-pairs ReduceScatter algorithm, which is not used in NCCL.
-In this algorithm, all GPUs send their data at $\frac{i}{N}$ offset to GPU $i$, the $i^{th}$ GPU does elementwise reduction of the received data, and stores the reduced data.
-To store the received data, the implementation allocates a scratch buffer per GPU.
-The kernel’s main loop performs reduction in batches of the scratch size (line 518).
-On each GPU, the loop first put $\frac{1}{N}$ of data to all other GPUs’ scratch buffer and then signals the completion of transfer (line 910).
-Then a GPU waits for the data to arrive from other GPUs and reduce it with its part of input data (line 1315).
-Finally, a barrier among all GPUs is required to ensure scratch is not overwritten for next main loop offset (line 18).
-Section 5 shows that this algorithm works better than the ring algorithm for smaller data sizes while the ring algorithm works better for large data sizes.
-3.2.3. Advantages of Communication Primitives
-1global allPairsRS(count, gpus, channels[gpus])
-2 sz = channel.scratchSz/gpus.num
-3 count = count/gpus.num
-4
-5 for (off = 0; off < count; off += sz)
-6 //Send 1/Nth data to each GPU
-7 for (g = 0; g < gpus.num; g++)
-8 idx = off + g * count;
-9 channels[g].put(idx, sz*g, sz)
-10 channels[g].signal()
-11
-12 //Reduce every pair of GPU
-13 for (g = 0; g < gpus.num-1; g++)
-14 channels[g].wait()
-15 reduce(in + off, channels[g].scratch)
-16
-17 //barrier on all gpus
-18 multiDeviceBarrier();
-Figure 5. All-pairs ReduceScatter kernel in MSCCL++. Channels are initialized with source as input and destination as scratch buffer.
-The MSCCL++ Primitive API itself does not introduce any novel optimization techniques – instead, it offers new straightforward abstractions that enable quick optimization of GPU communication.
-The key advantages come from exposing the primitive functionalities as a user interface.
-First, it allows application developers to easily and quickly optimize GPU communication algorithms for their own infrastructure and workloads.
-This is valuable in practice because the library developers may not be aware of the specific workloads and hardware configurations of their users.
-Second, the library developers can quickly deliver new hardware support by implementing the same primitive API for the new hardware.
-Third, since the primitive API is close to the hardware, it is more flexible to extend or adopt it for new hardware features.
-MSCCL++ carefully optimizes GPU communication in the DSL Executor and collective kernels leveraging some of the below features in the Primitive API.
-Users of the Primitive API can also implement these optimizations directly.
-Asynchronous Communication.
-MSCCL++ contains separate primitives for data transfer and synchronization, allowing asynchronous communication and batching synchronization of multiple data transfers.
-This abstraction has multiple benefits: (i) allows more algorithms to be specified than the synchronous communication in NCCL, (ii) enables faster collective communication performance by enabling algorithms to combine several synchronizations into one, (iii) frees up resources of peer GPUs for other meaningful work of the application, and (iv) enables better inter-kernel optimizations.
-Specialized GPU Kernels.
-By using channels for specific GPU interconnects, MSCCL++ allows users to specialize kernels for particular interconnects.
-Furthermore, MSCCL++ avoid extra copies in kernels, leading to a kernel with less code paths and zero register spills.
-For example, for eight A100 GPUs, the ring AllReduce of NCCL and MSCCL use 94 and 96 registers per thread, respectively, while an MSCCL++ version uses only 32 per thread.
-Due to these reasons, MSCCL++ kernels execute less instructions, avoid memory accesses due to register spills, and have better instruction cache hits.
-Interconnect Optimizations.
-NCCL only allows a single mode for data transfer over NVLink at a time.
-In contrast, by using different channel types in MSCCL++ (PortChannel or MemoryChannel), user can utilize both thread-copy and DMA-copy data transfer mode over NVLink.
-1class Channel
-2 protected:
-3 //Src/Dst Buffers set during initialization
-4 void* src, *dst;
-5 //Semaphore alloced on this GPU
-6 uint *semaphore; uint expectedVal;
-7 public:
-8 //Primitives
-9 void signal() ; void wait()
-10
-11class PortChannel : public Channel
-12 public:
-13 void put(ulong dstOff, ulong srcOff, ulong sz,
-14 uint tid, uint tids)
-15
-16template<Protocol protocol>
-17class MemoryChannel : public Channel
-18 public:
-19 void put(ulong dstOff, ulong srcOff, ulong sz,
-20 uint tid, uint tids, uint flag=0)
-21 template<typename T>
-22 T read(ulong off, uint tid, uint flag=0)
-23
-24 template<typename T>
-25 void write(ulong off, uint tid, T elem,
-26 uint flag=0)
-27
-28class SwitchChannel : public Channel
-29 public:
-30 void reduce(ulong dstOff, ulong srcOff,
-31 ulong sz, uint tid, uint tids)
-32 void broadcast(ulong dstOff, ulong srcOff,
-33 ulong sz, uint tid, uint tids)
-Figure 6. PortChannel, MemoryChannel, and SwitchChannel primitives in MSCCL++. Implementation of MemoryChannel primitives differ for the type of protocol: LL or HB.
-4. MSCCL++ Design and Implementation
-4.1. Initialization of MSCCL++
-In general, a MSCCL++ program runs as a multi-process on all nodes and GPUs within each node, so that each process runs one GPU.
-MSCCL++ provides host (CPU) side interfaces for initializing multiple processes, and the interfaces can be specialized by users for their application.
-The initialization involves setting up connections between GPUs, finding the topology of GPU connections, and constructing channels and buffers.
-The first step in initialization is to exchange metadata between host processes, which we call a Bootstrap.
-It consists of four virtual methods: (i) send to send metadata to another process, (ii) recv to receive metadata from a process, (iii) allGather to collect metadata of all processes, and (iv) barrier to synchronize all processes.
-MSCCL++ implements a default bootstrap methods using POSIX sockets, while it is straightforward to override with other popular implementations such as MPI or torch.distributed.
-The second step is to construct a Communicator that registers communication buffers (either user buffers or dedicated ones) and creates corresponding MSCCL++ channels between GPUs based on the underlying physical links.
-4.2. MSCCL++ Primitive API and Channels
-A MSCCL++ channel is a connection between two or more GPUs.
-PortChannel and MemoryChannel are peer-to-peer (i.e., between two GPUs), whereas SwitchChannel is for a group of (i.e., two or more) GPUs.
-Figure 6 shows three types of channels and its interface in MSCCL++.
-A channel is initialized with the source and destination buffers, and an integer allocated on the GPU memory that servers as a semaphore.
-4.2.1. PortChannel
-A PortChannel implements primitives when data transfer is done over ports connected to GPU memory, such as cudaMemcpy for intra-node DMA or ibv_post_send for RDMA.
-Since data transfer over a port currently requires the CPU to initiate the transfer, each channel creates its own CPU thread.
-The CPU thread reads data transfer or synchronization requests from a first-in-first-out request queue.
-The storage, head, and tail of a request queue are allocated using cudaMallocManaged as both CPU and GPU can access it.
-We now discuss the workflow of PortChannels using Figure 7.
-0 When the GPU calls a primitive, such as put, the first participating thread of the GPU pushes the request to the queue by writing at the head.
-Before writing to the queue, the GPU checks if the queue is filled, i.e., if the head value is more than the tail value.
-If the queue is filled then the GPU waits for the CPU to process atleast one request.
-1 Then the GPU increments the head to the next element.
-2 The CPU thread continuously reads the element at the tail to see if there is a request from the GPU, and when there is a request, the CPU reads the request, zeros out the current element, and increments the tail.
-3 Now the CPU thread will process the request.
-Below we explain the requests generated by primitives and how the CPU handle these requests.
 Figure 4. PortChannel workflow for IB from ⓪ when GPU 0 calls put primitive to ⑧ when GPU 1 receives the data.
@@
 6 7 The CPU processes this request by atomically incrementing the semaphore on the receiving GPU using functions, like ibv_atomic_add for IB.
-The wait primitive on the receiving GPU do not create a request for its CPU, therefore, the receiving CPU is idle, rather wait continuously looks for the semaphore to reach a expected value in a while loop.
+The wait primitive on the receiving GPU does not create a request for its CPU, therefore, the receiving CPU is idle, rather wait continuously looks for the semaphore to reach an expected value in a while loop.
 8 wait returns after the semaphore is incremented.
 Flush.
-Primitive flush pushes a flush request in the queue and the first thread of the GPU waits until the queue head is equal to or more than the queue tail.
+Primitive flush pushes a flush request in the queue and the first thread of the GPU waits until the queue tail reaches the position where the flush request is pushed to.
 The CPU processes the flush request by waiting until all the previous data transfer and synchronization requests have been completed.
@@
 After the flush request is complete, the GPU is free to re-write to the source buffer.
-4.2.2. MemoryChannel
-A MemoryChannel wraps data transfer methods that use thread-copy mechanism, i.e., directly use GPU threads for writing to peer GPUs memory.
+4.2. MemoryChannel
+A MemoryChannel wraps data transfer methods that use thread-copy mechanism, i.e., directly use GPU threads for writing to peer GPU’s memory.
 Protocols.
 The channel provides two protocols, which define data transfer and synchronization technique to tune between low-latency and high-bandwidth: HB protocol provides a high-bandwidth but high-latency protocol, thus, is suitable for larger sizes, and LL protocol provides low-latency but low-bandwidth, thus, is suitable for smaller sizes.
-Both protocols achieves these properties by synchronizing on the suitable granularity of data.
+Both protocols achieve these properties by synchronizing on the suitable granularity of data.
 (i) In the HB protocol, peer GPUs transfer a large chunk of data and synchronize this chunk once using wait and signal primitives.
@@
 However, since the receiving GPU needs to wait for the whole chunk before processing it, the HB protocol also has high latency.
-(ii) In LL protocol, the receiving GPU synchronizes on fewer chunks of data being transferred and process the chunk as soon as it is received, thus, achieving lower latency.
+(ii) In LL protocol, the receiving GPU synchronizes on fewer chunks of data being transferred and processes the chunk as soon as it is received, thus achieving lower latency.
 However, since the number of synchronization is proportional to the number of elements, this protocol provides lower-bandwidth.
@@
 The receiving GPU uses the read primitive, waits until the flag value at $N$ index of receiving buffer is set, and then reads and returns the N1N-1 elements.
-However, we cannot arbitrarily use any $N$ because GPUs follow a weak memory consistency model where writes to different memory locations by multiple threads can be performed in any order.
+However, we cannot arbitrarily use any $N$ because GPUs follow a weak memory consistency model in which writes to different memory locations by multiple threads can be performed in any order.
 Therefore, we restrict $N$ to number of elements written by a single instruction, i.e., 4, 8, and 16 bytes memory accesses instructions.
 The flag value is decided based on the algorithm, such that, all flag values are distinct.
-Implementation.
-A MemoryChannel is initialized with a protocol.
-A protocol defines the implementation of data transfer and synchronization primitives.
-We now describe the implementation of primitives based on the protocol as follows:
 Data Transfer.
-Primitive put reads data elements from source buffer and writes to destination buffer starting from different offsets.
+put is a collective-thread operations that reads data elements from source buffer and writes to destination buffer starting from different offsets using a set of threads.
 Similarly, the read primitive reads and returns data from the destination buffer, and the write primitive writes data including the flag for the LL protocol to the destination buffer.
 To maximize bandwidth for the HB protocol, both primitives use 16-bytes loads and stores.
-For the LL protocol, both primitives also takes a flag and uses 8-byte load and stores by default or user supplied vector length.
+For the LL protocol, both primitives also take a flag and use 8-byte loads and stores by default or user supplied vector length.
 Since both put primitives are called by multiple threads (all or first few threads of the kernel), we achieve maximum bandwidth.
@@
 The flush primitive is a no-op because after put returns, the source buffer can be reused, even though the write is still in progress.
-4.2.3. SwitchChannel
-A SwitchChannel provides primitives for performing collective operations among GPUs.
+4.3. SwitchChannel
+A SwitchChannel provides two primitives for performing collective operations among GPUs: reduce to add the corresponding elements of buffers residing on different GPUs and broadcast to send elements from a buffer on a GPU to all other GPUs.
 These operations usually require specialized hardware support.
-For example, NVIDIA NVLink 4.0 connects all H100 GPUs to a single NVSwitch and this NVSwitch can perform collective operations including reduce and broadcast.
-A SwitchChannel in MSCCL++ provides two primitives: reduce to add the corresponding elements of buffers residing on different GPUs and broadcast to send elements from a buffer on a GPU to all other GPUs.
-We now discuss the implementation of these primitives for NVSwitch, which provides in-network, switch-based aggregation and multicast capabilities using NVLink SHARP (NVLS) technology (NVIDIA, 2025c).
-However, we believe that these implementations can be generalized to other hardware in the future.
+For example, NVIDIA NVLink 4.0 connects all H100 GPUs to a single NVSwitch and this NVSwitch can perform in-network, switch-based aggregation and multicast using NVLink SHARP (NVLS) technology (NVIDIA, 2025a).
+We now discuss the implementation of these primitives for NVSwitch, but we believe that these implementations can be generalized to other hardware in the future.
 Reduce.
@@
 This primitive takes a source buffer as a local address on the local GPU and a destination buffer as a multimem address.
-The primitive goes through each element of the source buffer on the local GPU, reads the element into a register, and executes the multimem.st PTX instruction using the source register and the destination element’s multimem address.
-The multimem instruction sends the register value to the switch, which broadcasts and stores the value to all the virtual memory addresses pointed by the multimem address.
-4.3. MSCCL++ DSL API and Executor
-MSCCL++ DSL API is a Python-based language that allows describing a communication algorithm at a high level.
-It converts the algorithm description into a sequence of instructions that can be executed by the DSL Executor, which is a GPU kernel that reads and runs instructions back-to-back.
-MSCCL++ DSL is an extension of the MSCCLang (Cowan et al., 2023) that supports new instructions based on the MSCCL++ Primitive API and lifts a few restrictions of the language; e.g., we enable a single GPU thread block to access multiple GPUs at the same time.
-4.4. MSCCL++ Collective API and Kernels
-MSCCL++ Collective API is the simplest API that MSCCL++ offers, and it is the same as the popular NCCL API.
-We implement this API using kernels written in the MSCCL++ Primitive API.
-Users can either use these default kernels or plug-in their own algorithms written using the MSCCL++ DSL or Primitive APIs.
-We now briefly describe a few AllReduce algorithms included in the default kernels to: (i) provide ideas on low-level considerations for implementing high-performance GPU communication kernels, and (ii) to show how low-level interfaces can be useful for customization on different workloads.
+The primitive goes through each element of the source buffer on the local GPU, reads the element into a register, and executes a multimem.st PTX instruction using the source register and the destination element’s multimem address.
+The multimem instruction sends the register value to the switch, which broadcasts and stores the value to all virtual memory addresses pointed by the multimem address.
+4.4. Advantages of Primitive API
+Expert developers can use the Primitive API to build custom workload-specific communication algorithms in a portable manner with minimal performance overhead (see Table 1).
+At the same time, developers implementing these primitives can add support for emerging hardware and interconnects.
+By separating primitives for data transfer and synchronization, MSCCL++ allows asynchronous communication and batching synchronization of multiple data transfers.
+By using channels for specific GPU interconnects, MSCCL++ allows users to specialize kernels for particular interconnects.
+Furthermore, MSCCL++ avoids extra copies in kernels, leading to a kernel with less code paths and zero register spills.
+Since the Primitive API is an in-kernel interface for communication, we can co-optimize both computation and the communication kernels, leading to better application performance.
+One example is overlapping reduction in AllReduce with communication (see Section 5.2).
+Best Achievable MSCCL++
+NVLink Throughput (GB/s) 397.5 397.5
+NVLink Latency (ns) 822 829
+InfiniBand Throughput (GB/s) 48.94 48.94
+InfiniBand Latency ( µs) 3.76 4.89
+Table 1.
+MSCCL++ Primitive API’s peer-to-peer performance compared to the best achievable on our H100 environment (see Table 2).
+The best achievable performance is measured by nvbandwidth (NVIDIA, 2025c) and RDMA perftest (RDMA, 2025).
+MSCCL++’s NVLink throughput and latency are measured with PortChannel and MemoryChannel, respectively.
+MSCCL++’s InfiniBand throughput and latency are measured with PortChannel.
+5. MSCCL++DSL
+1class PortChannel:
+2 def __init__(self, dst_rank: int,
+3 src_rank: int): ...
+4 def put(self, dst_chunk: Buffer,
+5 src_chunk: Buffer, tb: int): ...
+6 def signal(self, tb: int): ...
+7 def wait(self, tb: int): ...
+8 def flush(self, tb: int): ...
+9
+10class MemoryChannel:
+11 def __init__(self, dst_rank: int,
+12 src_rank: int): ...
+13 def put(self, dst_chunk: Buffer,
+14 src_chunk: Buffer, tb: int,
+15 tb_group: ThreadBlockGroup): ...
+16 def put_packets(self, dst_chunk: Buffer,
+17 src_chunk: Buffer, tb: int,
+18 tb_group: ThreadBlockGroup): ...
+19 def reduce(self, dst_chunk: Buffer,
+20 src_chunk: Buffer, reduce_op: ReduceOp,
+21 tb: int, tb_group: ThreadBlockGroup): ...
+22 def signal(self, tb: int): ...
+23 def wait(self, tb: int): ...
+24
+25class SwitchChannel:
+26 def __init__(self, ranks: List[int]): ...
+27 def broadcast(self,
+28 dst_multimem_chunk: MultimemBuffer,
+29 src_chunk: Buffer, tb: int): ...
+30 def reduce(self, dst_chunk: Buffer,
+31 src_multimem_chunk: MultimemBuffer,
+32 reduce_op: ReduceOp, tb: int): ...
+Figure 5. Channel abstractions in MSCCL++ DSL (Python).
+To enable specifying custom communication algorithms, MSCCL++ supports a Python-based Domain-Specific Language (DSL) and lowers the algorithm into a sequence of operations that can be executed by the DSL Executor on top of the primitive abstractions described above.
+The primitives exposed at the DSL retain the zero-copy, one-sided, and asynchronous properties of the underlying primitive abstractions, thereby enabling users to specify optimized algorithms that overlap compute and communication.
+1def ringRS(portChannels, bufs, sz, N, tb=0):
+2 scr = [Buffer(rank, sz) for rank in range(N)]
+3 for rank in range(N):
+4 putChan = portChannels[(rank+1)%N]
+5 recvChan = portChannels[(rank+N-1)%N]
+6 src = bufs[rank]
+7 recv, dst = scr[rank], scr[(rank+1)%N]
+8
+9 def chunk(step):
+10 csz = sz / N
+11 offset = ((rank + N - step) % N) * csz
+12 return offset, offset+csz/2, offset+csz
+13
+14 for step in range(N):
+15 beg, mid, end = chunk(step)
+16 # (a) Put 1st half of chunk
+17 putChan.put(dst[beg:mid], src[beg:mid], tb)
+18 putChan.signal(tb)
+19 if step != 0:
+20 pbeg, pmid, pend = chunk(step-1)
+21 # Reduce prev 2nd half, overlaps (a)
+22 src[pmid:pend].reduce(recv[pmid:pend], tb)
+23 recvChan.wait(tb) # Wait for 1st half
+24 putChan.flush(tb) # Flush 1st half
+25 # (b) Put 2nd half of chunk
+26 putChan.put(dst[mid:end], src[mid:end], tb)
+27 putChan.signal(tb)
+28 # Reduce 1st half, overlaps (b)
+29 src[beg:mid].reduce(recv[beg:mid], tb)
+30 recvChan.wait(tb) # Wait for 2nd half
+31 putChan.flush(tb) # Flush 2nd half
+32 if step == N-1:
+33 src[mid:end].reduce(recv[mid:end], tb)
+Figure 6. A simplified example of a Ring ReduceScatter implementation using MSCCL++ DSL.
+5.1. DSL
+MSCCL++DSL provides a global view of all thread blocks on all GPUs or ranks in a Python-native language.
+As shown in Figure 5, the DSL exposes PortChannel, MemoryChannel, and SwitchChannel abstractions in this global view.
+Consequently, unlike the Primitive API (Section 4), users specify the ranks involved in each channel and for each operation in the channel, specify the thread block on the channel’s source rank that must execute it.
+Some operations on the MemoryChannel also accept an optional ThreadBlockGroup that enables a group of thread blocks to collectively perform that operation.
+The Channel classes include a few fused primitives (omitted in Figure 5 for brevity) that can reduce the overhead of API calls.
+For example, MemoryChannel’s reduce_put reduces two local buffers and puts the result to a remote destination using a temporary register to avoid a memory round-trip for the intermediate data.
+In addition, the DSL provides a separate Buffer abstraction and the Channels are not tied to a pair of buffers, so users specify the chunks of the source and destination buffer in each data transfer operations.
+Since the DSL is Python-native, the chunks are specified as slices of Buffer.
+Buffers are associated with a rank and allow computation like reduce, which is performed on its rank.
+The DSL also provides convenience classes to create input and output buffers based on the communication pattern (e.g., AllReduce, AllGather, ReduceScatter), device-wide synchronization primitives, and multi-device synchronizations primitives; synchronization within a thread block is handled automatically by the DSL.
+Users can control chunk sizes, number of queue pairs (QPs), etc.
+via the DSL.
+For instance, if users want two QPs, they can use two parallel PortChannels.
+If users want to offload the communication to CPU, they can use PortChannel rather than MemoryChannel.
+Thus, the MSCCL++ DSL enables users to construct complex data movement and synchronization workflows without writing low-level CUDA code.
+5.2. Example: Overlapped Ring ReduceScatter
+Figure 6 demonstrates how the DSL can implement an efficient Ring ReduceScatter algorithm that overlaps computation and communication.
+In the classic Ring algorithm, $N$ GPUs are organized in a ring, the input buffer is divided into $N$ chunks, and each chunk circulates around the ring to be reduced.
+At each step, a GPU receives a chunk, reduces it locally, and then forwards it to the next GPU, with all steps occurring synchronously.
+To enable overlap, Figure 6 further splits each chunk into two halves.
+This allows the reduction of the first half (Line 29) to overlap with the communication of the second half (Line 26), and the reduction of the second half (Line 22) to overlap with the communication of the first half (Line 17).
+In Figure 6, only thread block 0 on each GPU does this computation and communication.
+The user can experiment with more parallelism or pipelining by using more thread blocks on each GPU and/or using more chunks per GPU.
+The MSCCL++ DSL supports replicating program instances to increase parallelism and improve performance while automatically handling mapping of the program to thread blocks and channels.
+5.3. Lowering and Optimizations
+The program written in MSCCL++ DSL is lowered to an execution plan, which consists of information from the program such as the collective, channel types, data transfer protocol, memory buffers to register for each thread block, semaphores to setup, and the sequence of operations (which may include loops) to run in each thread block of every rank.
+Buffer sizes and rank counts (N in Figure 6) are provided by the user at lowering time.
+During lowering, the program goes through data dependence analysis to add appropriate synchronizations, as well as optimization passes such as operation fusion.
+This enables the DSL to achieve performance comparable to hand-tuned CUDA implementations.
+Data Dependence Analysis.
+MSCCL++ DSL automatically tracks data dependencies at the chunk level within each thread block by maintaining the last writer and active readers for each memory slot.
+When operations have data dependencies, the lowered program includes necessary synchronization points to ensure correct execution order.
+For example, in Figure 6, it introduces a thread block synchronization on thread block 0 before the reduce operation (Line 29).
+The DSL also detects redundant synchronizations and removes them.
+Specifically, if a lowered program will be calling multiple thread block synchronizations back-to-back, the redundancies will be removed, retaining only one of them.
+Operation Fusion.
+MSCCL++ DSL also maintains a directed acyclic graph (DAG) to track data dependencies and usage patterns of operations in a thread block at the chunk level.
+When two or more operations meet the fusion criteria, such as contiguous chunk access, no intervening dependencies, and compatible resource requirements, the DSL merges them into a single operation function.
+This fusion strategy reduces memory traffic and avoids unnecessary synchronization, resulting in faster execution.
+For example, a DSL code using MemoryChannel:
+src.reduce(data,tb); memChan.put(dst,src,tb);is captured into a fused operation:
+memChan.reduce_put(dst,src,data,tb);where reduce_put reduces two local buffers and puts the result to a remote destination using a temporary register to avoid a memory round-trip for the intermediate data.
 Env. Name GPU Intra-node Link Network
@@
 Table 2. List of environments used for evaluation.
+5.4. Executor
+Given an execution plan and the input and output buffers, the DSL Executor initializes different channels, setting up connections and semaphores, and registering memory.
+Finally, the executor runs an execution kernel that interprets the list of operations in the execution plan.
+The execution kernel is a single GPU kernel implementation that runs any given execution plan.
+The execution kernel inlines calls to primitive MSCCL++ operations such as put, signal, wait, and flush.
+The execution plan can also contain fused operations, such as ReduceSend explained in Section 5.3.
+6. MSCCL++Collectives Library
+MSCCL++Collective API is the simplest API that MSCCL++ offers, and it is the same as the popular NCCL or RCCL API.
+We implement this API using kernels written in the MSCCL++ DSL.
+Users can either use these default kernels or plug-in their own algorithms written using the MSCCL++ DSL or Primitive APIs.
+We now briefly describe a few AllReduce algorithms included in the default kernels to: (i) provide ideas on low-level considerations for implementing high-performance GPU communication kernels, and (ii) to show how low-level interfaces can be useful for customization on different workloads.
 1.
@@
 As we use the 1PA algorithm only for very small message collectives within a single node, we implement only a single version of 1PA that uses the LL protocol with MemoryChannel.
-MSCCL++ can implement it much more efficiently than other libraries by removing synchronizations.
+MSCCL++ can implement it much more efficiently than other libraries by relaxing synchronizations and concurrently transferring data to multiple devices before waiting on any of them.
 2.
@@
 As another example, we can let a single thread group read data from multiple other GPUs at the same time.
-This allows efficient data reduction compared with other libraries that read data from different GPUs one-by-one, which synchronizes each reduction step.
+This allows efficient data reduction compared with other libraries that read data from different GPUs one-by-one, which synchronizes at each reduction step.
 3.
+Two-phase Ring (2PR).
+Similar to 2PA, 2PR has two phases: the first phase does ReduceScatter in a ring and the second phase does AllGather in the same ring.
+Unlike NCCL, we can use PortChannel (DMA-copy) even within a node and we overlap the reduction with the DMA-copy.
+Figure 6 shows a simplified kernel for ReduceScatter.
+We extend such pipelining to overlap reduction with communication across both phases.
+Our 2PR implementation using PortChannel shows the best throughput among all implementations for intra-node AllReduce with large message sizes.
+4.
 Two-phase Hierarchical (2PH).
@@
 Unlike the first variant, to utilize the link bandwidth efficiently, the number of data chunks is the same as the number of GPUs.
-4.5. Portability and Implementation Effort
-MSCCL++ is designed to be portable.
-Code using MSCCL++ collectives, algorithms using the MSCCL++ DSL, and kernels using the MSCCL++ primitives are portable as long as the underlying MSCCL++ Primitive API supports that hardware.
-Our design makes it easy to add support for new hardware.
-As an example, MSCCL++ initially supported only NVIDIA GPUs; to add support for AMD MI300x GPUs, our implementation took only 7 weeks for one developer: 3 weeks for basic AMD GPU support and 4 weeks to develop new AllReduce algorithms that outperform RCCL/MSCCL for message sizes of 1KB - 1GB.
-It is especially remarkable that the AMD-specific code in MSCCL++ (excluding makefiles and algorithms) is only less than 10 lines of code, while RCCL, the hard-forked NCCL for AMD GPUs, is substantially diverged from NCCL.
-MSCCL++ could achieve this because the low-level API of AMD GPUs (i.e., HIP) is almost the same as that of NVIDIA GPUs (i.e., CUDA), and the MSCCL++ Primitive API is only a shallow abstraction on top of the low-level API.
-This significantly reduces the effort to maintain and develop MSCCL++.
-The MSCCL++ design makes it easy to extend the Primitive API to support new hardware features.
-For example, NVIDIA NVLink 4.0 introduced support for multimem instructions for switch-based aggregation and multicast.
-Since this feature introduced a very different concept from previous interfaces (i.e., PortChannel or MemoryChannel), we developed the SwitchChannel interface to support this feature.
-The development took only 8 weeks for two developers, including learning the basic usage of this feature, abstracting the feature as a new type of channel (i.e., SwitchChannel), and finally developing a new AllReduce algorithm using SwitchChannel that outperforms NCCL/MSCCL by more than $2.2\times$ in average for message sizes of 1KB - 1GB.
 7. Evaluation
 Environments.
-All presented numbers are collected from one of the environments listed in Table 1.
-Each node is equipped with 8 GPUs per node (either A100, H100, or MI300X), intra-node links between GPUs (either NVLink or Infinity Fabric (a.k.a.
-xGMI)), and inter-node IB links (one NIC per GPU).
+Table 2 lists the environments used in the evaluation.
+Each node is equipped with 8 GPUs per node (either A100, H100, or MI300x) with intra-node links between GPUs (either NVLink or Infinity Fabric (a.k.a.
+xGMI)) and inter-node IB links (one NIC per GPU).
 All NICs are connected to a single IB networking switch.
@@
 For brevity, the rest of this paper refers to each environment by the name of GPU.
-Baselines.
-We compare MSCCL++ with the state-of-the-art collective communication libraries, including NCCL 2.26.2 (NVIDIA, 2025a), RCCL 2.20.5 (AMD, 2025), [^5] As RCCL partly adopts MSCCL, to prevent confusion, we refer to MSCCL-disabled RCCL.
-and MSCCL 2.23 (Microsoft, 2025a).
-MSCCL borrows the stack implementation of NCCL/RCCL and supports custom algorithms.
-Therefore, MSCCL performs exactly the same as NCCL/RCCL if they use the same algorithm.
-5.1. Collective Communication
-To compare the performance of collective communication, we implement the AllReduce algorithms described in Section 4.4 using the MSCCL++ DSL API.
+Baselines. We compare MSCCL++ with the state-of-the-art collective communication libraries, including NCCL 2.26.2 (NVIDIA, 2025d), RCCL 2.20.5 (AMD, 2025), and MSCCL 2.23 (Microsoft, 2025a).
+Figure 7. AllReduce, A100-40G. MnNg means M nodes and N GPUs. X-axis is message size (at the bottom of the figure).
+7.1. Collective Communication Results
+To compare the performance of collective communication, we implement the AllReduce algorithms described in Section 6 using the MSCCL++ DSL.
 We also implement various AllGather algorithms in a similar manner, and we omit the implementation details.
-We present the best number among all implementations for each message size and environment.
-All NCCL, RCCL, and MSCCL numbers are fine-tuned for each environment and message size by adjusting their environment variables, such as the number of channels (affects the number of threads), chunk size (affects the size of data to be transferred at once), type of algorithm (such as ring, tree, or NVLS (NVIDIA, 2025c)), the topology (XML file that describes the intra-node link topology of GPUs), etc.
+We present the best number among all implementations for each message size and environment [^3] We did not do sophisticated autotuning beyond picking the best configuration for a given platform and input size ranges (akin to MSCCLang (Cowan et al., 2023)) by offline profiling.
+Our experiments measure the performance of the resulting system.
+Orthogonal techniques from TACCL (Shah et al., 2023) and TE-CCL (Liu et al., 2024) can be used for automation, which is an interesting future direction..
+All NCCL, RCCL, and MSCCL numbers are fine-tuned for each environment and message size by adjusting their environment variables, such as the number of channels (affects the number of threads), chunk size (affects the size of data to be transferred at once), type of algorithm (such as ring, tree, or NVLS (NVIDIA, 2025a)), the topology (XML file that describes the intra-node link topology of GPUs), etc.
 For MSCCL, we use the fastest algorithm for each message size (Microsoft, 2025b).
-We use NCCL’s user buffer registration API (i.e., ncclMemAlloc(NVIDIA, 2025f) and the CUDA/HIP Graph APIs for best performance.
-For visibility, we separate the range of message sizes into small (up to 1MB, presented latency) and large (1MB and above, presented algorithm bandwidth (AlgoBW) [^6] Defined as the message size divided by the latency.) in the figures.
+We use NCCL’s user buffer registration API (i.e., ncclMemAlloc(NVIDIA, 2025h) and the CUDA/HIP Graph APIs for best performance.
+For clarity of exposition, we separate message sizes into small (up to 1MB, presented as latency) and large (1MB and above, presented as algorithm bandwidth (AlgoBW)) in the figures.
 The small message sizes represent inference scenarios (such as LLM token sampling a.k.a.
@@
 For small messages, MSCCL++ outperforms NCCL and MSCCL by up to 5.4x and 2.3x, respectively, and for large messages, MSCCL++ outperforms NCCL and MSCCL by up to 1.8x and 1.4x, respectively.
-For AllGather, we observe a few cases where MSCCL slightly outperforms MSCCL++ by up to 8.0%, which is introduced by a performance bug of the DSL API, not by our algorithm implementation.
-For AllReduce, MSCCL++ is always the fastest.
-Figure 8. AllReduce, A100-40G. MnNg means M nodes and N GPUs. X-axis is message size (at the bottom of the figure).
+MSCCL++ is faster than other baselines across the board for AllReduce.
+For AllGather, MSCCL outperforms MSCCL++ by 8.0% in one scenario.
+This is due to the performance overhead of our DSL API that we are able to reduce if we implement directly on top of MSCCL++’s Primitive API.
+Figure 8. AllGather, A100-40G. MnNg means M nodes and N GPUs. X-axis is message size (at the bottom of the figure).
 Gain Breakdown.
-The performance differences can come from both high-level differences in algorithms (i.e., design of data flow routes and pipelines) and low-level optimizations (i.e., how to transfer data, synchronize, and manage buffers).
-The former benefit is already delivered by MSCCL that enables the customization of algorithms over the NCCL communication stack.
-Therefore, the key benefits of MSCCL++ lie in the latter by exposing the primitive functionalities to users.
-Since MSCCL and MSCCL++ use the same high-level algorithm for all cases in Figure 8 and 9, the performance gap between the two comes purely from the benefits of MSCCL++.
+The performance gain of MSCCL++ comes from both algorithmic improvements, fine-grained overlapping of compute and communication, and the ability to exploit low-level optimization capabilities provided by the Primitive API.
 NCCL vs MSCCL.
@@
 For very large message sizes $\geq$ 256MB in multi-node cases, MSCCL achieves higher bandwidth than NCCL because it uses a hierarchical algorithm (2PH), which is more bandwidth-efficient.
-Figure 9. AllGather, A100-40G. MnNg means M nodes and N GPUs. X-axis is message size (at the bottom of the figure).
 […]
@@
 We have implemented both the DSL and Primitive versions up to 2-node experiments to verify that their performance is similar, while 4-node experiments are done only using the DSL.
-However, as the DSL introduces higher level instructions than the Primitive API, DSL versions perform 3% worse than the Primitive versions on average, and are up to 18% worse in one corner case.
-Despite limitations, MSCCL++ DSL is still useful for quick prototyping of collective communication and easy-to-understand algorithm description.
-Collective vs Primitive. For AI inference workloads, we implement the best algorithms in our collective kernels using the MSCCL++ Primitive API, so there is no performance difference.
-Figure 10. Speedup for decodes with tensor parallelism = 8
-5.2. LLM Inference
-[…]
+As the DSL introduces a runtime interpreter executor, DSL versions perform 3% worse than the Primitive versions on average, and are up to 18% worse in one corner case.
+Despite performance overheads, MSCCL++ DSL is useful for quick prototyping of collective communication with easy-to-understand algorithm description.
 7.2. Cross-Hardware Support
-To show the effectiveness of MSCCL++ on different hardware platforms, Figure 11 and 12 compare the single-node AllReduce performance on H100 and MI300x.
+To show the effectiveness of MSCCL++ on different hardware platforms, Figure 9 and 10 compare the single-node AllReduce performance on H100 and MI300x.
 […]
@@
 This means, for best link utilization, we need to copy data to all peers at the same time as much as possible, unlike for NVIDIA GPUs where we can copy data to each peer one-by-one back-to-back.
-Using the MSCCL++ DSL, this is as easy as changing the order of two nested for loops, one that walks through data elements and the other that walks through different peers.
+Using the MSCCL++ DSL, this is as easy as changing the order of two nested for loops, one that walks through data elements and the other that walks through peers.
+Figure 11. Speedup of MSCCL++ over NCCL for Llama3-70b decodes with tensor parallelism = 8.
+7.3. LLM Inference
+[…]
+vLLM also has a custom hand-written AllReduce kernel for a single node.
+For different message sizes, MSCCL++’s AllReduce performs similar or up to $3\times$ faster than the custom kernel, with a geomean improvement of $1.4\times$.
+For end-to-end decode latency, MSCCL++ performs similar or up to $1.11\times$ (on average $1.04\times$) faster than inference with the custom AllReduce kernel.
+[…]
+Figure 12. DeepSeek-V3 decode throughput using SGLang on two H100 nodes (16 GPUs) with tensor parallelism (1024 input tokens and 1024 output tokens per batch).
+Figure 13. DeepEP on two H100 nodes (16 GPUs) with expert parallelism (DeepSeek-V3 settings: hidden size 7168, top-k 8, 256 total experts, FP8 dispatch and BF16 combine).
+Expert Parallelism with DeepEP.
+DeepEP (Zhao et al., 2025) is an implementation of expert parallelism for Mixture-of-Experts (MoE) layers, which is proposed by DeepSeek-V3 (DeepSeek-AI, 2024) for both training and inference.
+The original code leverages a hardware feature called InfiniBand GPUDirect Async (IBGDA) (Markthub et al., 2022) (featured by NVSHMEM (NVIDIA, 2025e)) that implements an InfiniBand networking stack inside GPU, so that GPU can directly communicate with NIC bypassing CPU.
+However, this feature is only available on NVIDIA GPUs combined with Mellanox NICs at the time of writing.
+To make it portable to other environments, we integrate MSCCL++ into DeepEP to replace the IBGDA-based networking functions into PortChannel API calls, which uses the RDMA networking stack running on CPU instead (see Figure 4).
+[^4] IBGDA can be supported by the MSCCL++ PortChannel API in the future.
+It took a couple of weeks for one developer to add this support, but most of that time was spent in understanding DeepEP code.
+We evaluated DeepEP using its official benchmark and setup.
+IBGDA was enabled for all experiments, which was leveraged by NVSHMEM, but not by MSCCL++.
+Figure 13 demonstrates that the two implementations exhibit no noticeable performance difference.
+Unlike NCCL, MSCCL++ is flexible enough to implement DeepEP-like applications efficiently.
+The IBGDA stack implementation in DeepEP is less portable and hard to maintain, while MSCCL++’s PortChannel simplifies this code drastically while being portable and performant.
+7.4. Portability and Implementation Effort
+[…]
+The MSCCL++ design makes it easy to extend the Primitive API to support new hardware features.
+For example, NVIDIA NVLink 4.0 introduced support for multimem instructions for switch-based aggregation and multicast.
+Since this feature introduced a very different concept from previous interfaces (i.e., PortChannel or MemoryChannel), we developed the SwitchChannel interface to support this feature.
+The development took only 8 weeks for two developers, including learning the basic usage of this feature, abstracting the feature as a new type of channel (i.e., SwitchChannel), and finally developing a new AllReduce algorithm using SwitchChannel that outperforms NCCL/MSCCL by more than $2.2\times$ on average for message sizes of 1KB - 1GB.
 8. Related Work
-NVIDIA Collective Communications Library (NCCL) (NVIDIA, 2025a) and AMD ROCm Collective Communications Library (AMD, 2025) (RCCL) are vendor supplied libraries for NVIDIA and AMD GPUs respectively.
-Cowan et al.
-(Cowan et al., 2022) improved over these libraries by allowing execution of collective algorithms specialized for a size and topology.
-In this paper, we showed that primitives provided by MSCCL++ performs better than the primitives of these libraries.
-The primitive interface of MSCCL++ has similarities with some other works in terms of being a GPU-side API for communication.
-For example, the NVIDIA OpenSHMEM (NVSHMEM) (NVIDIA, 2025b) is a parallel work that provides primitive functions such as nvshmemx_putmem_warp as well as a few collective functions such as nvshmemx_broadcastmem_warp.
-However, we could not find any implementation where NVSHMEM outperforms NCCL (or MSCCL++) for collective communication, and quantitative comparison is difficult as NVSHMEM is not open source.
+GPU Communication Primitives: NVSHMEM (NVIDIA, 2025e) is an extension of the OpenSHMEM (OpenSHMEM, 2025) API for NVIDIA GPUs for providing a shared-memory abstraction.
+In contrast, MSCCL++ provides performance-preserving portable abstractions and higher-level DSL for developing efficient communication collective algorithms.
+Concurrent to our work, NVSHMEM recently introduced support for inter-node communication (after MSCCL++ was open sourced).
+MSCCL++ provides different interconnect abstractions, namely PortChannel, MemoryChannel, and SwitchChannel, to enable optimized implementations.
+As an example, NVSHMEM’s put always uses thread-copy when it is feasible, so users cannot choose DMA-copy on NVLink.
+Similarly, NVSHMEM exposes raw multimem capabilities by exposing underlying memory pointers and users need to implement their kernels using PTX instructions.
+With an eye towards portability, SwitchChannel abstracts low-level details with a uniform synchronization mechanisms.
+Finally, NVSHMEM does not provide an API like MemoryChannel’s LL protocol.
+It may be possible to implement such an API using NVSHMEM, but that would add burden on users.
 ARK (Hwang et al., 2023b) is another work that proposes a GPU-side control plane for communication, but it is implemented as a monolithic end-to-end ML system rather than a standalone communication library.
-UCX, an open-source framework developed by Shamis et al.
-(Shamis et al., 2015) and libraries implementing the MPI standard (Thakur et al., 2005) offer a comprehensive set of APIs for high-performance computing (HPC) environments.
-In contrast, MSCCL++ is more specific to GPU acceleration.
-SCCL (Cai et al., 2021), TACCL (Shah et al., 2023), and TE-CCL (Liu et al., 2024) aim to accelerate GPU collective communication by synthesizing efficient data transfer algorithms for performing collective communication, while still using the vendor-provided primitives underneath.
-The MSCCL++ primitives can be used with the generated algorithms to capture their benefit.
-Finally, LLM inference frameworks like vLLM (Kwon et al., 2023) and TensorRT-LLM (NVIDIA, 2025e) implement custom AllReduce kernels for achieving high performance communication.
-However, the custom implementations are not general-purpose and often only limited to single-node collective communication.
-MSCCL++, on the other hand, enables communication across multiple types of interconnects and has equivalent AllReduce performance to vLLM’s custom single-node AllReduce.
-7. Conclusion
+DSL: Cowan et al.
+(Cowan et al., 2022) introduced MSCCLang, a DSL to specify collective communication algorithms specialized for a message size and network topology.
+MSCCLang is built on top of NCCL and RCCL send-recv abstraction, so is limited to two-sided, synchronous communication.
+Our MSCCL++ DSL provides a one-sided, asynchronous API that enables expressing more efficient communication algorithms.
+Due to this, MSCCL++ is faster than MSCCL, which is built using MSCCLang DSL, as shown in our evaluation.
+Communication Libraries: NVIDIA Collective Communications Library (NCCL) (NVIDIA, 2025d) and AMD ROCm Collective Communications Library (AMD, 2025) (RCCL) are vendor supplied libraries for NVIDIA and AMD GPUs respectively.
+We showed that MSCCL++ Collectives perform better than both for AI inference workloads.
+RCCL has since adopted MSCCL++.
+Communication Algorithms: SCCL (Cai et al., 2021), TACCL (Shah et al., 2023), and TE-CCL (Liu et al., 2024) aim to accelerate GPU collective communication by synthesizing efficient data transfer algorithms using the NCCL and RCCL send-recv abstraction.
+Their techniques are orthogonal and their performance may improve if they generate algorithms using the MSCCL++ Primitives API.
+LLM Inference Frameworks: LLM inference frameworks like SGLang (SGLang, 2025), DeepEP (Zhao et al., 2025), vLLM (Kwon et al., 2023), and TensorRT-LLM (NVIDIA, 2025g) implement custom kernels for achieving high performance communication.
+However, the custom implementations are not general-purpose and often only limited to specific interconnects like NVLink or xGMI.
+MSCCL++ not only performs similar to their custom implementations on their corresponding interconnects, but also supports efficient communication across multiple types of interconnects.
+SGLang has since adopted MSCCL++.
+9. Conclusion and Future Work
 MSCCL++is a novel GPU communication stack designed for high-performance AI applications.
-By exposing the primitive communication functionalities as straightforward user interfaces, MSCCL++ enable fine-grained optimizations for GPU experts, while also providing higher-level interfaces for quick optimizations.
+By exposing the primitive communication functionalities as straightforward user interfaces, MSCCL++ enables fine-grained optimizations for GPU experts, while also providing higher-level interfaces for quick optimizations.
 Moreover, such a design can reduce the overall development and optimization effort for GPU communication, and accelerates adoption of fast evolving hardware technologies.
-By implementing collective communication using the proposed MSCCL++ interfaces, we can achieve up to $5.4\times$ speedup for standalone collectives, and up to $15\%$ speedup for end-to-end AI inference.
+By implementing collective communication using the proposed MSCCL++ interfaces, we can achieve up to $5.4\times$ speedup for standalone collectives, and up to $1.38\times$ speedup for end-to-end AI inference.
+MSCCL++ could be used for accelerating AI training workloads too, but we leave that for future work.
+MSCCL++ could also evolve into a general I/O abstraction for AI accelerators in the future; e.g., by supporting non-GPU accelerators and file system I/O transports like BlitzScale (Zhang et al., 2025).
+Appendix A Artifact Appendix
+A.1. Abstract
+The artifact includes implementations of the three APIs described in this paper, namely the Primitive API (C++/CUDA) for low-level optimization of GPU communication, the DSL API (Python) that is implemented over the Primitive API and enables high-level description of custom collective communication algorithms, and the Collective API (C++) that implements the NCCL/RCCL API as-is over the Primitive API and the DSL API.
+The Collective API can be used to reproduce collective communication results in Section 7.1 and 7.2.
+A.2. Artifact check-list (meta-information)
+  • Program: The latest version of nccl-tests (or rccl-tests). Publicly available and not included.
+  • Run-time environment: Linux; Ubuntu 20.04 or later is recommended.
+Docker is recommended, otherwise need manual installations of dependencies such as CUDA (or ROCm), Python3, MPI, and optionally OFED for multi-node experiments.
+  • Hardware: 8 datacenter-grade GPUs. Optionally RDMA NICs for multi-node experiments.
+  • Execution: Sole user recommended for best performance.
+  • Metrics: Latency and bandwidth are reported.
+  • Output: Console output with reported metrics.
+  • Experiments: Step-by-step guide in documentation.
+  • How much disk space required (approximately)?: ¡1GB for the source code and a build; 10GB for a Docker image with all dependencies.
+  • How much time is needed to prepare workflow (approximately)?: 10 minutes
+  • How much time is needed to complete experiments (approximately)?: 1 hour
+  • Publicly available?: Yes
+  • Code licenses (if publicly available)?: MIT License
+  • Archived (provide DOI)?: 10.5281/zenodo.18334822
+A.3. Description
+A.3.1. How to access
+The source code is publicly available on GitHub at https://github.com/microsoft/mscclpp. The repository is sized less than 1MB, and a build would be a few hundreds of MB.
+A.3.2. Hardware dependencies
+At least 2 GPUs are needed for functionality tests, and 8 or more are needed for reproducing numbers in Section 7.1 and 7.2.
+Datacenter-grade GPUs with high-bandwidth interconnects (NVLink or Infinity Fabric) are recommended (see Table 2).
+Optionally for multi-node experiments, RDMA NICs with InfiniBand or RoCE configurations are required.
+A.3.3. Software dependencies
+CUDA (or ROCm), Python3, and MPI are required for artifact evaluation. Optionally OFED is required for multi-node experiments. See details in the ”Prerequisites” section in docs/quickstart.md.
+A.4. Installation
+  1. (1)
+    Download and run a provided Docker image. This step can be skipped if the host environment is preferred. See ”Docker Images” section in docs/quickstart.md.
+  2. (2)
+    Build from the source.
+See the step-by-step guide in ”Install from Source” section in docs/quickstart.md.
+If the environment is not configured for multi-node execution, we recommend to disable InfiniBand/RoCE features via a CMake option -DMSCCLPP_USE_IB=OFF.
+  3. (3)
+    Build and run unit_tests and mp_unit_tests for functionality testing. See the step-by-step guide in ”Unit Tests” section in docs/quickstart.md.
+A.5. Evaluation and expected results
+For evaluation, install the latest version of nccl-tests (for NVIDIA GPUs) or rccl-tests (for AMD GPUs), which are publicly available at https://github.com/NVIDIA/nccl-tests and https://github.com/ROCm/rccl-tests, respectively.
+MPI should be enabled during their compilation (set environment variable MPI=1 during make).
+After that, follow the ”NCCL/RCCL Benchmark over MSCCL++” section in docs/quickstart.md to run nccl-tests or rccl-tests using the MSCCL++ Collective API.
+The testing binaries will print latency and bandwidth numbers on the console, similar to what presented in Figure 7, 8, 9, and 10, depending on the GPU type and the number of GPUs.
+Please refer to the README.md of nccl-tests for detailed usages.
+A.6. Experiment customization
+The implementation of MSCCL++ Collective API can be customized using two different interfaces.
+  • Using the Primitive API. Users can write their own GPU communication kernel using the MSCCL++ Primitive API and let the Collective API invoke the custom kernel. We provide a simple example in
+    docs/guide/customized-algorithm-with-nccl-api.md.
+  • Using the DSL API.
+Users can leverage our Python DSL to describe a collective communication algorithm that can be lowered and executed by the Collective API.
+We provide a simple example in docs/dsl_quick_start.md.
+A.7. Notes
+All documentation is also publicly available at https://microsoft.github.io/mscclpp.
+A.8. Methodology
+Submission, reviewing and badging methodology: